### v2

# Via P. Giuria 1, 10125 Torino, Italy

Abstract
We initiate the lightcone bootstrap analysis of multipoint correlators in a defect conformal field theory. The setup we consider is the three-point function of two bulk and one defect operator. Requiring consistency of the crossing equation in the lightcone limit, we find constraints on the defect spectrum at large transverse spin. Specifically, to reproduce the exchange of the leading-twist operator in the bulk channel we find two new twist-accumulating families of defect operators at large transverse spin and we compute their defect CFT data in this limit.
1 Introduction and discussion
The central idea of the conformal bootstrap program is to constrain the dynamics of conformal field theories (CFTs) using symmetries and internal consistency. After the breakthrough of Rattazzi:2008pe , this program was revived, leading to important numerical and analytical results (see Rychkov:2023wsd ; Bissi:2022mrs and references therein). One of the early achievements of the analytic bootstrap is due to Komargodski:2012ek ; Fitzpatrick:2012yx , where it was found that a generic CFT must contain an infinite tower of double-twist operators, whose dimension approaches the generalized free field (GFF) dimension at large spin. The main idea leading to this result was to exploit the crossing equation for the four-point function of identical operators in Lorentzian signature and consider its behaviour in the lightcone limit, i.e. when two operators become lightlike separated. Several important developments followed, either trying to make the statement mathematically rigorous Qiao:2017xif ; Kravchuk:2020scc ; Kravchuk:2021kwe ; Pal:2022vqc ; vanRees:2024xkb or extending and exploiting the initial result Fitzpatrick:2014vua ; Alday:2015eya ; Kaviraj:2015cxa ; Li:2015itl ; Komargodski:2016gci ; Alday:2016njk ; Simmons-Duffin:2016wlq ; Kusuki:2021gpt ; Numasawa:2022cni ; Pal:2023cgk . In particular, the Lorentzian inversion formula for four-point correlators encoded the initial observation of Komargodski:2012ek ; Fitzpatrick:2012yx in a general framework which allows to extract the CFT data from the discontinuity of the correlator, thus providing an efficient computational method for all those theories that admit a perturbative expansion and some control on the perturbative spectrum Caron-Huot:2017vep .
The line of development we are interested in is the extension of the lightcone bootstrap program to multipoint correlators Antunes:2021kmm ; Kaviraj:2022wbw ; Harris:2024nmr . Extending the bootstrap program to higher-point correlation functions provides access to a larger set of constraints and to a new set of exchanged local operators Rosenhaus:2018zqn ; Parikh:2019ygo ; Fortin:2019dnq ; Goncalves:2019znr ; Parikh:2019dvm ; Pal:2020dqf ; Fortin:2020bfq ; Bercini:2020msp ; Buric:2020dyz ; Fortin:2020zxw ; Poland:2021xjs ; Buric:2021ywo ; Buric:2021ttm ; Buric:2021kgy ; Fortin:2022grf ; Poland:2023vpn ; Fortin:2023xqq ; Antunes:2023kyz ; Poland:2023bny ; Harris:2025cxx ; Antunes:2025vvl . In particular, the lightcone bootstrap for four identical scalar operators gives access to the large-spin spectrum of double-twist operators and their OPE coefficients with the external scalars. On the other hand, considering five- and six-point functions significantly enlarges the set of CFT data, providing access to OPE coefficients of spinning operators as well as the spectrum and OPE coefficients of triple-twist operators. The drawback is that the kinematics of higher-point functions is significantly more intricate and this makes it harder to derive closed-form expressions for the CFT data.
In this paper, we analyze a setup that enlarges the set of available CFT data with a milder complication of the kinematical structure. To achieve this, we consider a CFT in the presence of a conformal defect, i.e. an extended excitation which preserves conformal invariance along its profile. Starting with the seminal work of Billo:2016cpy , the application of bootstrap techniques to conformal defects has evolved into an independent research line, the defect bootstrap program. If one is interested in the interaction between bulk and defect degrees of freedom, the first observable with a non-trivial crossing equation is the two-point function of bulk operators. The lack of positivity in the bulk channel does not allow for a straightforward application of numerical bootstrap methods (see however Meineri:2025qwz ; Lanzetta:2025xfw for recent progress), but it is certainly possible to apply analytic bootstrap techniques. In particular, Lorentzian inversion formulae Lemos:2017vnx ; Liendo:2019jpu and dispersion relations Bianchi:2022ppi ; Barrat:2022psm have been derived for both OPE channels and their simplicity allowed for the computation of the correlator in situations where one can expand in a small parameter having some control on the spectrum of exchanged operators Bissi:2018mcq ; Dey:2020jlc ; Barrat:2021yvp ; Liendo:2016ymz ; Gimenez-Grau:2019hez ; Gimenez-Grau:2021wiv ; Bianchi:2022sbz ; Gimenez-Grau:2022ebb ; Bianchi:2023gkk ; Chen:2023yvw ; Chen:2024orp .
Figure 1: Examples of multipoint defect correlators with their respective number of conformal cross-ratios.
Here, we are interested in extending Lorentzian bootstrap methods to multipoint defect correlators. One of the nice features of defect correlators is that the set of available configurations is extremely rich. One option, of course, is to consider several defect operators (see, for instance, Artico:2024wut for a perturbative application in superconformal field theories), but we would like to probe also the interaction with the bulk. Some higher-point functions with at least one bulk operator are shown in Figure 1, where we also indicate the number of conformal cross-ratios associated to each correlator. The number of cross-ratios could be used as a measure of the complexity of the kinematical structure. To compare with the homogeneous case, we should remember that the four-point function of local operators is a function of two cross-ratios (exactly like the two-point function of bulk operators with a defect), while the five- and six-point functions can be expressed in terms of five and nine cross-ratios, respectively. As we see in Figure 1, for defect correlators the difficulty grows at a slower rate as we can construct intermediate observables which are just slightly more complicated than the bulk two-point function. Among those, the simplest examples are the BDDD and the BBD correlators [^1] One may wonder why the BDD correlator is not included in this list. It turns out that the BDD correlator is even simpler than the BB one as it is characterized by a single cross-ratio and there is no crossing equation (see however Artico:2024wnt ; Bartlett-Tisdall:2025iqx for recent analyses of this setup).. The former however, does not probe new bulk-to-defect CFT data and for this reason in this paper we focus on the latter.
Summary of the results
The main goal of this work is to extract information about the spectrum and OPE coefficients of defect operators with large spin by requiring consistency of the crossing equation in the lightcone limit. The class of defect operators that can be constrained using the bulk two-point function of two identical scalars $\expectationvalue{\Phi\Phi}$ are the so-called transverse derivative operators and they take the schematic form $\Phi]_{m,s}=\partial_{\perp}^{s}\square_{\perp}^{m}\Phi$, where the operator $\Phi$ is taken on the defect and the derivatives act on the orthogonal directions Lemos:2017vnx . Using consistency of the crossing equation in the lightcone limit one can prove that any defect conformal field theory must contain this family of operators in their defect spectrum and their half-twist at infinite spin is $\hat{h}=\frac{\hat{\Delta}-s}{2}=\frac{\Delta_{\Phi}}{2}+m$.
Figure 2: A schematic representation of the crossing equation for the bulk-bulk-defect correlator. The red curly line indicates the $\bar{z}\to 1$ lightcone limit we are considering, where the two bulk operators become lightlike separated. The defect operator $\hat{\phi}$ acts as a spectator in both OPE channel decompositions.
Here we consider the correlator $\big\langle\Phi\Phi\hat{\phi}\big\rangle$, where we added a defect operator $\hat{\phi}$. This correlator admits two different OPE expansions, which are schematically represented in Figure 2. We study the resulting crossing equation in the limit where the two bulk operators become lightlike separated as shown in Figure 2 through a red curly line. In the bulk channel, the exchanged operator with lowest twist provides the leading contribution in the lightcone limit. A crucial difference compared to the BB case is that the bulk identity operator is not exchanged as it does not couple with a generic defect operator $\hat{\phi}$ (unless $\hat{\phi}$ is the defect identity of course). Therefore, to understand which operators dominate the bulk expansion at low twist we need some insight into the twist spectrum of the bulk theory. Focusing on unitary theories, we know that the unitarity bounds for the bulk theory impose $h\geq\frac{d-2}{4}$ for scalars and $h\geq\frac{d-2}{2}$ for spinning operators (remember that $h=\frac{\Delta-J}{2}$ is the half twist, so for the stress tensor $h=\frac{d-2}{2}$ and for a scalar $h=\frac{\Delta}{2}$). In this work we exclude higher-spin conserved currents as they are expected to be absent in an interacting CFT for $d>2$ Maldacena:2011jn . Therefore, there are no spinning operators saturating the unitarity bound for $J>2$. The leading-twist contribution is given either by the stress tensor operator or by a light scalar operator (with $\frac{d-2}{2}\leq\Delta\leq d-2$) [^2] Strictly speaking, one should consider also a spin-one conserved current, but the two-point function of a bulk current with a defect scalar operator vanishes for generic codimension. It could be non-vanishing for specific cases where parity-odd contributions must be included, see e.g. Bianchi:2019sxz ; Bachas:2025brp . In this work we analyze both these options and we study the implications for the defect spectrum at large spin.
In the defect channel, the exchanged operators $\hat{\mathcal{O}}_{\hat{h}_{1},s}$ and $\hat{\mathcal{O}}_{\hat{h}_{2},s}$ must have the same spin, which is taken large to reproduce the bulk channel, but they have different transverse twists $\hat{h}_{1}$ and $\hat{h}_{2}$. In particular, to reproduce the lowest-twist contribution in the bulk channel we need to introduce two new families of exchanged defect operators:
  1. 1.
    “Double-twist” operators [^3] The language is slightly misleading, but we consider as “single twist” operators the transverse derivative operators with $\hat{h}=\frac{\Delta_{\Phi}}{2}+m$ and we label “double-twist” and “triple-twist” those with $\hat{h}=\frac{\Delta_{\Phi}}{2}+\frac{\hat{\Delta}}{2}+m_{1}+m_{2}$ and $\hat{h}=\frac{\Delta_{\Phi}}{2}+\hat{\Delta}+m_{1}+m_{2}$ respectively. of the schematic form
    $\hat{\phi}\Phi]_{m_{1},m_{2},s}=\hat{\phi}\Box^{m_{1}}_{\parallel}\Box_{\perp}^{m_{2}}\partial^{s}_{\perp}\Phi\,$, (1)
    with half-twist
    $\hat{h}=\frac{\Delta_{\Phi}}{2}+\frac{\hat{\Delta}_{\hat{\phi}}}{2}+m_{1}+m_{2}$ (2)
    at infinite spin
  2. 2.
    “Triple-twist” operators of the schematic form
    $\hat{\phi}\hat{\phi}\Phi]_{m_{1},m_{2},s}=\hat{\phi}^{2}\Box^{m_{1}}_{\parallel}\Box_{\perp}^{m_{2}}\partial^{s}_{\perp}\Phi\,$, (3)
    with half-twist
    $\hat{h}=\frac{\Delta_{\Phi}}{2}+\hat{\Delta}_{\hat{\phi}}+m_{1}+m_{2}$ (4)
    at infinite spin.
Requiring consistency of the crossing equation in the lightcone limit we can find an expression for the large-spin behavior of the bulk-to-defect couplings $b_{\Phi[\hat{\phi}\Phi]}$ and $b_{\Phi[\hat{\phi}\hat{\phi}\Phi]}$ in terms of the bulk OPE data, i.e. the product of a bulk OPE coefficient $\lambda_{\Phi\Phi\mathcal{O}_{\star}}$ and a bulk-to-defect coupling $b_{\mathcal{O}_{\star}\hat{\phi}}$, where $\mathcal{O}_{\star}$ is the leading-twist exchanged operator, i.e. the stress tensor or a light scalar operator. Finding closed-form expressions for these couplings at arbitrary $m_{1}$ and $m_{2}$ is not easy, but we provide a relation that allows to extract these OPE coefficients recursively in $m_{1}$ and $m_{2}$. Furthermore, thanks to the simplicity of the defect setup, we can obtain closed-form expressions for the cases $m_{1}=0$ or $m_{2}=0$ and arbitrary values of the other index. Here we show only the result for a scalar operator. More general expressions for arbitrary spin can be found in Section 3. For a scalar bulk exchange of dimension $\Delta=2h_{\star}$ we find that the leading behavior for the defect CFT data at large spin is
$b_{\Phi\widehat{\mathcal{O}}_{1}}b_{\Phi\widehat{\mathcal{O}}_{2}}\lambda_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}\hat{\phi}}\simeq C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}\frac{2^{s}s^{2h_{\Phi}-h_{\star}-1}}{\Gamma(2h_{\Phi}-h_{\star})}.$ (5)
Notice that the crossing equation is only sensitive to the product $b_{\Phi\widehat{\mathcal{O}}_{1}}b_{\Phi\widehat{\mathcal{O}}_{2}}\lambda_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}\hat{\phi}}$ as expected. However, to reproduce the bulk expansion we observe two possibilities
  1. 1.
    $\widehat{\mathcal{O}}_{1}=[\Phi]_{m_{2},s}$ and $\widehat{\mathcal{O}}_{2}=[\hat{\phi}\Phi]_{m_{1},m_{2},s}$. In this case we can compute the product $b_{\Phi[\Phi]}b_{\Phi[\hat{\phi}\Phi]}\lambda_{[\Phi][\hat{\phi}\Phi]\hat{\phi}}$, but the large spin behavior of $b_{\Phi[\Phi]}$ can be extracted studying the $\expectationvalue{\Phi\Phi}$ correlator and the OPE coefficient $\lambda_{[\Phi][\hat{\phi}\Phi]\hat{\phi}}$ appears in the defect four-point function $\expectationvalue{\hat{\phi}[\Phi]\hat{\phi}[\Phi]}$. Therefore the only new (not accessible through lower points correlators) piece of defect CFT data is $b_{\Phi[\hat{\phi}\Phi]}$. The coefficient $C_{[\hat{\phi}\Phi]_{m_{1},m_{2},s}[\Phi]_{m_{2},s}}$ can be computed in a closed form either for $m_{1}=0$ or for $m_{2}=0$. The results are
    $C_{[\hat{\phi}\Phi]_{0,m,s}[\Phi]_{m,s}}=\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b_{\mathcal{O}_{\star}\hat{\phi}}\frac{\Gamma(2h_{\star})\Gamma(m+h_{\star})\Gamma(1-\hat{h}_{\hat{\phi}})\Gamma(\hat{h}_{\hat{\phi}})\Gamma(m-\hat{h}_{\hat{\phi}}+h_{\star})}{\Gamma(h_{\star}){}^{2}\Gamma(m+1)\Gamma(h_{\star}-\hat{h}_{\hat{\phi}})\Gamma(\hat{h}_{\hat{\phi}}+h_{\star})\Gamma(m-\hat{h}_{\hat{\phi}}+1)}$ (6)
    $\times\ _{3}F_{2}(-m,\hat{h}_{\hat{\phi}}-m,-\frac{d}{2}+2h_{\Phi}-h_{\star}+1;-m-h_{\star}+1,-m+\hat{h}_{\hat{\phi}}-h_{\star}+1;1)$
    $C_{[\hat{\phi}\Phi]_{m,0,s}[\Phi]_{0,s}}=\lambda_{\Phi\Phi\mathcal{O}_{\star}}b_{\mathcal{O}_{\star}\hat{\phi}}\frac{2\Gamma(2h_{\star})\Gamma(\hat{h}_{\hat{\phi}}+1)\Gamma(2(m+\hat{h}_{\hat{\phi}}))\,}{\Gamma(h_{\star})\Gamma(m+1)\Gamma(\hat{h}_{\hat{\phi}}+h_{\star})\Gamma(m+2\hat{h}_{\hat{\phi}}+1)}$
    $\times\ _{3}F_{2}(-m,\hat{h}_{\hat{\phi}}+1,2\hat{h}_{\hat{\phi}};m+2\hat{h}_{\hat{\phi}}+1,\hat{h}_{\hat{\phi}}+h_{\star};1)$
  2. 2.
    $\widehat{\mathcal{O}}_{1}=[\hat{\phi}\Phi]_{0,m_{2},s}$ and $\widehat{\mathcal{O}}_{2}=[\hat{\phi}\hat{\phi}\Phi]_{m_{1},m_{2},s}$. In this case we have access to the product $b_{\Phi[\hat{\phi}\Phi]}b_{\Phi[\hat{\phi}\hat{\phi}\Phi]}\lambda_{[\hat{\phi}\Phi][\hat{\phi}\hat{\phi}\Phi]\hat{\phi}}$ and the new piece of defect CFT data is $b_{\Phi[\hat{\phi}\hat{\phi}\Phi]}$. We can compute the coefficient $C_{[\hat{\phi}\hat{\phi}\Phi]_{0,m_{2},s}[\hat{\phi}\Phi]_{m_{1},m_{2},s}}$ in a closed form for $m_{1}=0$
    $\!C_{[\hat{\phi}\hat{\phi}\Phi]_{0,m,s}[\hat{\phi}\Phi]_{0,m,s}}\!\!=\!\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b_{\mathcal{O}_{\star}\hat{\phi}}\frac{\Gamma(2h_{\star})\Gamma(m+h_{\star})\Gamma(1+\hat{h}_{\hat{\phi}})\Gamma(-\hat{h}_{\hat{\phi}})\Gamma(m+\hat{h}_{\hat{\phi}}+h_{\star})}{\Gamma(h_{\star}){}^{2}\Gamma(m+1)\Gamma(h_{\star}+\hat{h}_{\hat{\phi}})\Gamma(-\hat{h}_{\hat{\phi}}+h_{\star})\Gamma(m+\hat{h}_{\hat{\phi}}+1)}$ (7)
    $\times\ _{3}F_{2}(-m,-\hat{h}_{\hat{\phi}}-m,-\frac{d}{2}+2h_{\Phi}-h_{\star}+1;-m-h_{\star}+1,-m-\hat{h}_{\hat{\phi}}-h_{\star}+1;1)$
These results are very general and they hold for any unitary interacting CFT. As an application, we analyzed the specific case of the BBD three-point function of two chiral primary operators in the bulk and one tilt operator on the defect in $\mathcal{N}=4$ SYM. In that case, the spectrum of exchanged operators in the bulk channel is well known and the leading-twist contribution comes from operators in the stress tensor supermultiplet. When specified to that case, our results significantly simplify and we collect them in Section 4.
2 Review: bulk two-point correlators and lightcone bootstrap
To set the stage for higher-point applications, let us review how the lightcone bootstrap works for bulk two-point functions Lemos:2017vnx . We consider a flat $p$-dimensional defect, and we denote the codimension of the defect with q=dpq=d-p. We will always assume that $q>1$  [^4] For $q=1$ the kinematics is different and there is no notion of transverse spin.. We work in the embedding space formalism, which we review in Appendix A. The bulk two-point function of identical scalar operators reads
$\expectationvalue{\Phi_{\Delta_{\Phi}}(P_{1})\Phi_{\Delta_{\Phi}}(P_{2})}=\frac{f(\xi,\cos\varphi)}{\left(P_{1}\!\circ\!P_{1}\right)^{\frac{\Delta_{\Phi}}{2}}\left(P_{2}\!\circ\!P_{2}\right)^{\frac{\Delta_{\Phi}}{2}}}$ (8)
with the cross ratios
$\xi=-\frac{2P_{1}\cdot P_{2}}{\left(P_{1}\!\circ\!P_{1}\right)^{\frac{1}{2}}\left(P_{2}\!\circ\!P_{2}\right)^{\frac{1}{2}}}\,,\qquad\cos\varphi=\frac{P_{1}\!\circ\!P_{2}}{\left(P_{1}\!\circ\!P_{1}\right)^{\frac{1}{2}}\left(P_{2}\!\circ\!P_{2}\right)^{\frac{1}{2}}}\,.$ (9)
To understand the geometry of the problem in Lorentzian signature it is convenient to consider the conformal frame depicted in Figure 3.
Figure 3: The role of the $z,\bar{z}$ cross ratios can be easily understood in the conformal frame where the two bulk operators lie on a Minkowski plane orthogonal to the defect $\mathcal{D}$. The $z=0$ and $\bar{z}=0$ lines represent the defect lightcone, while the $z=1$ and $\bar{z}=1$ lines correspond to the two bulk operators being lightlike separated.
The convenient cross ratios in that frame are $z$ and $\bar{z}$ given by
$\xi=\frac{(1-z)(1-\bar{z})}{(z\bar{z})^{\frac{1}{2}}}\,,\qquad\cos\varphi=\frac{z+\bar{z}}{2(z\bar{z})^{\frac{1}{2}}}\,.$ (10)
The function $f(z,\bar{z})$ admits two different block expansions as shown in Figure 4.
Figure 4: A schematic representation of the crossing equation for the bulk two-point function. The red curly line indicates the $\bar{z}\to 1$ lightcone limit we are considering, where the two bulk operators become lightlike separated.
In the defect channel, we can express the correlator as
$f(z,\bar{z})=\sum_{\hat{h},s}b^{2}_{\hat{h},s}\hat{f}_{\hat{h},s}(z,\bar{z})\,$, (11)
where the defect conformal blocks $\hat{f}_{\hat{h},s}(z,\bar{z})$ are known explicitly Billo:2016cpy ; Lemos:2017vnx
$\hat{f}_{\hat{h},s}(z,\bar{z})=z^{\hat{h}}\bar{z}^{\hat{h}+s}{}_{2}F_{1}\left(-s,\frac{q}{2}-1,2-\frac{q}{2}-s,\frac{z}{\bar{z}}\right)\,{}_{2}F_{1}\left(2\hat{h}+s,\frac{p}{2},2\hat{h}+s+1-\frac{p}{2},z\bar{z}\right)\,.$ (12)
In (11) we are summing over the transverse spin $s$ and the half transverse twist $\hat{h}=\frac{\hat{\Delta}-s}{2}$ of the defect exchanged operator. The coefficients $b_{\hat{h},s}$ are bulk-to-defect couplings determined by
$\expectationvalue{\Phi_{\Delta_{\Phi}}(P_{1})\widehat{\mathcal{O}}_{\hat{h},s}(P_{2},W)}=\frac{b_{\hat{h},s}(P_{1}\!\circ\!W)^{s}}{(P_{1}\!\bullet\!P_{2})^{2\hat{h}+s}(P_{1}\!\circ\!P_{1})^{\frac{\Delta_{\Phi}}{2}-\hat{h}}}\,$. (13)
In the bulk channel, the correlator can be expanded as
$f(z,\bar{z})=\left(\frac{(1-z)(1-\bar{z})}{(z\bar{z})^{\frac{1}{2}}}\right)^{-{\Delta}_{\Phi}}\sum_{\Delta,J}\lambda_{\Phi\Phi\mathcal{O}}\,a_{\mathcal{O}}f_{\Delta,J}(z,\bar{z})\,$, (14)
where we sum over the dimension $\Delta$ and the spin $J$ of the bulk exchanged operator. The bulk conformal blocks $f_{\Delta,J}(z,\bar{z})$ can be expressed as solutions to a Casimir differential equation as we review in Appendix C. The coefficients $\lambda_{\Phi\Phi\mathcal{O}}$ are bulk three-point functions, while the coefficients $a_{\mathcal{O}}$ appear in the one-point function
$\expectationvalue{\mathcal{O}_{\Delta,J}(P,Z)}=a_{\mathcal{O}}\frac{\left(\frac{(P\circ Z)^{2}}{P\circ P}-Z\!\circ\!Z\right)^{\frac{J}{2}}}{\left(P\!\circ\!P\right)^{\Delta+J}}\,.$ (15)
We are interested in a regime where
$1-\bar{z}\ll z<1\,$, (16)
which corresponds to a double lightcone limit approached in a specific direction as depicted in Figure 3. Specifically, we will work at leading order for $\bar{z}\to 1$ and at all orders for $z\to 0$. At this point, it is important to mention that the procedure we follow in this paper, both for this review and for the higher-point generalization, is not completely rigorous as it was addressed in a series of papers Qiao:2017xif ; Kravchuk:2020scc ; Kravchuk:2021kwe ; Pal:2022vqc ; vanRees:2024xkb for the homogeneous case. In this first work on the subject, we neglect these subtleties and we assume that all the nice features of the OPE expansion in the double lightcone limit that worked for the homogeneous case keep working also for the defect case.
In the limit $\bar{z}\to 1$ the bulk expansion converges and we can use (14) to evaluate $f(z,\bar{z})$ for $\bar{z}\to 1$. In particular, the bulk block in the lightcone limit reads
$f_{\Delta,J}(z,\bar{z})=(1-\bar{z})^{\frac{\Delta-J}{2}}\left(\tilde{f}_{\Delta+J}(z)+O(1-\bar{z})\right)\,$, (17)
where the function $\tilde{f}_{\Delta+J}(z)$ is given in (85). This implies that the leading contribution is determined by exchanged operators with the lowest twist. When considering two identical external operators the lowest-twist exchanged operator is clearly the identity. In that case $\Delta=J=0$ and the correlator is simply given by the prefactor in (14). The crossing equation states that this contribution must be reproduced in the defect channel and this imposes constraints on the large-spin spectrum of defect operators. In particular, one can reproduce the $(1-\bar{z})^{-\Delta_{\Phi}}$ factor by considering a scaling limit such that $s\to\infty$ and $\bar{z}\to 1$ with $s(1-\bar{z})$ kept fixed. Expanding the defect blocks in this limit (see Appendix D for details), we get
$\hat{f}_{\hat{h},s}(z,\bar{z})\sim z^{\hat{h}}(1-z)^{1-\frac{d}{2}}e^{-s(1-\bar{z})}\,.$ (18)
To reproduce the contribution of the bulk identity order by order in $z\to 0$ we need to ask that the large-spin spectrum of the theory contains an infinite tower of operators with transverse twist
$\hat{h}=\frac{\Delta_{\Phi}}{2}+m+O(s^{-\alpha})\,,\qquad s\to\infty\,$, (19)
for $m\in\mathbb{N}$ and $\alpha>0$. As discussed in Lemos:2017vnx , these operators are single-twist defect operators of the schematic form $\partial_{\perp}^{s}\square_{\perp}^{m}\Phi$. They appear naturally in GFF and their leading role at large spin is the defect counterpart of the emergence of GFF in the large-spin spectrum of homogeneous CFTs Fitzpatrick:2012yx ; Komargodski:2012ek . To reproduce the prefactor in (14), one also needs to impose that the defect OPE coefficients of these single twist operators $b^{2}_{m,s}$ behave as a power law for $s\to\infty$
$b^{2}_{m,s}=s^{\Delta_{\Phi}-1}\left(\frac{1}{\Gamma(\Delta_{\Phi})}\left(\begin{array}[]{c}m-\frac{d}{2}+\Delta_{\Phi}\\ m\end{array}\right)+O(s^{-\beta})\right)\,$, (20)
for $\beta>0$. Inserting (20) and (18) into the defect block expansion (11), we can replace the sum over spins with an integral dominated by the large-spin region such that
$f(z,\bar{z})\sim\sum_{m=0}^{\infty}\left(\begin{array}[]{c}m-\frac{d}{2}+\Delta_{\Phi}\\ m\end{array}\right)\frac{z^{\frac{\Delta_{\Phi}}{2}+m}(1-z)^{1-\frac{d}{2}}}{\Gamma(\Delta_{\Phi})}\int_{\Lambda}^{\infty}e^{-s(1-\bar{z})}s^{\Delta_{\Phi}-1}ds\,$, (21)
for some cut-off $\Lambda$, whose value is unimportant in the $\bar{z}\to 1$ limit. The integral, at leading order for $\bar{z}\to 1$, reproduces the $(1-\bar{z})^{-\Delta_{\Phi}}$ factor in (14) yielding
$f(z,\bar{z})\sim(1-\bar{z})^{-\Delta_{\Phi}}(1-z)^{-\Delta_{\Phi}}z^{\frac{\Delta_{\Phi}}{2}}\,.$ (22)
This result agrees with the prefactor in (14) for $\bar{z}\to 1$. In principle, one could use the same strategy to reproduce the twist-subleading contributions in the bulk expansion (14). This constrains the large-spin corrections to the classical dimension (19) and to the OPE coefficients (20) and a systematic large-spin expansion can be established. We refer to Lemos:2017vnx for further details on this expansion, which can be understood more rigorously in the context of the Lorentzian inversion formula. For the the higher-point generalization, which presents a richer and more complicated kinematical structure, we will concentrate on the leading-twist contribution. As we will highlight in the next Section, the leading contribution is already highly non-trivial because the bulk-bulk-defect correlator does not include the identity in the bulk channel.
3 Multipoint defect correlators
In the previous section, we reviewed how the crossing equation for the bulk two-point correlator in a lightcone limit determines the asymptotics of bulk-to-defect OPE coefficients $b^{2}_{m,s}$ at large spin. Our aim now is to study the simplest extension of these constraints in the setting of higher-point correlation functions. The scalar higher-point correlators with simplest kinematics were summarized in Figure 1, where we also represented the number of cross-ratios associated with each case. As we argued there, the simplest non-trivial case which admits a crossing equation involving both bulk expansions and defect expansions is the Bulk-Bulk-Defect (BBD) correlator, which will be the main focus of the rest of this work.
3.1 The Bulk-Bulk-Defect bootstrap setup
A BBD correlator can be expressed as follows
$\expectationvalue{\Phi(P_{1})\Phi(P_{2})\hat{\phi}(P_{3})}=\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\!(P_{1},P_{2},P_{3})\,\mathcal{G}(z,\bar{z},x)\,$, (23)
in terms of a conventional prefactor,
$\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}=\frac{(-2P_{1}\bullet P_{2})^{\hat{\Delta}_{\hat{\phi}}/2}}{(-2P_{1}\bullet P_{3})^{\hat{\Delta}_{\hat{\phi}}/2}(-2P_{2}\bullet P_{3})^{\hat{\Delta}_{\hat{\phi}}/2}(P_{1}\circ P_{1})^{\Delta_{\Phi}/2}(P_{2}\circ P_{2})^{\Delta_{\Phi}/2}}\,$, (24)
and a function of three cross-ratios $\mathcal{G}(z,\bar{z},x)$. Two of these cross-ratios, $z$ and $\bar{z}$, are identical to the Bulk-Bulk cross-ratios defined in (10), while the third one, $x$, can be defined via
$\frac{x^{2}+z\bar{z}}{(1+x^{2})(1-z)(1-\bar{z})}=-\frac{\left(P_{1}\!\circ\!P_{1}\right)\left(P_{2}\!\bullet\!P_{3}\right)}{2\left(P_{1}\cdot P_{2}\right)\left(P_{1}\!\bullet\!P_{3}\right)}\,.$ (25)
These three degrees of freedom can be easily visualized in the conformal frame of Figure 5
Figure 5: The role of the $z,\bar{z}$ cross ratios is the same as for the case of the bulk two-point function. The two bulk operators lie on a Minkowski plane orthogonal to the defect $\mathcal{D}$. The defect operator $\hat{\phi}$ is located along the defect at a distance $x$ from the transverse plane.
where $x$ is the distance of the defect field insertion $\hat{\phi}$ from the $(z,\bar{z})$-plane where the two bulk fields $\Phi$ are located.
As before, we can expand the correlator in two different OPE channels as represented in Figure 2. The bulk-channel expansion has the form
$\mathcal{G}(z,\bar{z},x)=\sum_{\Delta,J,n}\lambda_{\Phi\Phi\mathcal{O}}\,b_{\mathcal{O}\hat{\phi}}^{(n)}\,f_{\Delta,J,n}\,(z,\bar{z},x)$ (26)
where $\Delta\in\mathds{R}_{+}$ and $J\in 2\mathbb{N}$ are the labels of the exchanged bulk primaries in the $\Phi\times\Phi$ OPE, while $n=0,\dots,J/2$ labels the possible tensor structures of the bulk-to-defect two-point function [^5] We remind the reader that our embedding space conventions, including the definition of the $C_{i}^{AI}$, are gathered in Appendix A.
$\expectationvalue{\!\mathcal{O}_{\Delta,J}(P_{1},Z_{1})\hat{\phi}(P_{2})\!}=\frac{\left(4C_{1}^{AI}P^{2}_{A}P^{1}_{I}\right)^{J}}{\left(-2P_{1}\!\bullet\!P_{2}\right)^{\hat{\Delta}_{\hat{\phi}}+J}\!\left(P_{1}\!\circ\!P_{1}\right)^{\frac{\Delta-\hat{\Delta}_{\hat{\phi}}+J}{2}}}\sum_{n}b_{\mathcal{O}\hat{\phi}}^{(n)}\!\left(\!\frac{C_{1}^{AI}C^{1}_{AI}\left(P_{1}\!\bullet\!P_{2}\right)^{2}}{2\left(C_{1}^{AI}P^{2}_{A}P^{1}_{I}\right)^{2}}\!\right)^{\frac{J}{2}-n}\hskip-5.0pt.$ (27)
Note that here we are considering general contributions for a generic $p$-dimensional defect. We are therefore not considering special cases where odd spins and parity-odd tensor structures can contribute, for which the expressions in (27) would only constitute a part of the answer.
The defect channel, instead, can be expressed as
$\mathcal{G}(z,\bar{z},x)=\sum_{\hat{h}_{1},\hat{h}_{2},s}b_{\Phi\widehat{\mathcal{O}}_{1}}b_{\Phi\widehat{\mathcal{O}}_{2}}\lambda_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}\hat{\phi}}\,\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}(z,\bar{z},x)\,.$ (28)
Here, the labels are the half transverse twists $\hat{h}_{1}$, $\hat{h}_{2}$ and the transverse spin $s_{1}=s_{2}\equiv s$ of the two operators $\widehat{\mathcal{O}}_{1}$, $\widehat{\mathcal{O}}_{2}$.
We aim to study the crossing equation
$\sum_{\Delta,J,n}\lambda_{\Phi\Phi\mathcal{O}}\,b_{\mathcal{O}\hat{\phi}}^{(n)}\,f_{\Delta,J,n}(z,\bar{z},x)=\sum_{\hat{h}_{1},\hat{h}_{2},s}b_{\Phi\widehat{\mathcal{O}}_{1}}b_{\Phi\widehat{\mathcal{O}}_{2}}\lambda_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}\hat{\phi}}\,\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}(z,\bar{z},x)$ (29)
in the limit $\bar{z}\to 1$, where the two bulk fields become null separated. This is a natural limit to consider in the bulk channel as it corresponds to the Lorentzian OPE limit, in which the expansion is dominated by the operators with lowest twist $2h=\Delta-J$.
The leading-twist exchange
Contrary to the bulk-bulk case, the identity exchange is not allowed in the bulk channel of the BBD correlator because the one-point function $\langle\hat{\phi}(P_{3})\rangle$ is zero (or, analogously, because $b_{\mathds{1}\hat{\phi}}=0$). This means that the first contribution in the $\bar{z}\to 1$ lightcone limit is theory-dependent and thus already non-trivial. This leading contribution corresponds to the allowed exchanged operators with the lowest possible twist above the identity $h_{\star}>0$. We will work under the assumption that there is only one such operator with lowest twist $h_{\star}$, which we dub the leading-twist operator $\mathcal{O}_{\star}$. This assumption is not very restrictive, as the case of a finite number of leading-twist operators with the same twist is simply covered by running our analysis term-by-term and recombining all contributions in the end. Due to the presence of an infinite family of operators approaching the spectrum of double-twist operators $\Phi\Phi]_{0,J}$ at large spin $J$, our assumption automatically imposes the bound $h_{\star}<2h_{\Phi}$. In an interacting unitary theory, the leading-twist operator typically corresponds to the stress tensor $T^{\mu\nu}$ with $2h_{T}=d-2$ and $J_{T}=2$, to a conserved current $\mathcal{J}^{\mu}$ with $2h_{\mathcal{J}}=d-2$ and $J_{\mathcal{J}}=1$, or to some scalar operator $\varphi$ with $\frac{d-2}{4}\leq h_{\varphi}\leq\frac{d-2}{2}$. The case of a conserved current will not be relevant for our analysis, as it cannot be exchanged in the OPE of two identical fields due to its spin being odd (see however footnote 2). Our goal is then to explore what constraints the presence of such a leading-twist operator implies in the defect channel expansion.
The bulk-channel expansion
To work out the bulk-channel block $f_{\Delta,J,n}(z,\bar{z},x)$ in the $\bar{z}\to 1$ lightcone limit, we use the integral expression for the leading term in the Lorentzian OPE Ferrara:1971zy ; Ferrara:1971tq ; Ferrara:1972cq
[…]
with $\tilde{P}=P_{2}-t(P_{2}-P_{1})$, $\tilde{Z}=P_{2}-P_{1}$, with $\mathrm{B}$ representing the Euler beta function, and where we dropped subleading terms in $P_{12}$. Assuming the presence of a single leading-twist operator $\mathcal{O}_{\star}$, the sum over $\mathcal{O}$ collapses to just one term. Plugging this expansion in the $\expectationvalue{\Phi(P_{1})\Phi(P_{2})\hat{\phi}(P_{3})}$ correlator leads to an integral of bulk-to-defect correlators (27), which can be evaluated in terms of a Lauricella $F_{D}$ function of three variables (see Appendix C for a more detailed derivation, including integral and series representations for this function). Making sure of extracting the prefactor (24) from this expression, we then get an analytic expression for the bulk-channel expansion in the $\bar{z}\to 1$ lightcone limit
$\mathcal{G}(z,\bar{z},x)\stackrel{{\scriptstyle\bar{z}\to 1}}{{\simeq}}\\ \sum_{n}\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}(1-\bar{z})^{h_{\star}-2h_{\Phi}}(1-z)^{\bar{h}_{\star}-2h_{\Phi}}z^{-\bar{h}_{\star}+h_{\Phi}+\hat{h}_{\hat{\phi}}}(x^{2}-z)^{2n}(x^{2}+z)^{-2n-\hat{h}_{\hat{\phi}}}\\ \times\left(\frac{1+x^{2}}{1+z}\right)^{\hat{h}_{\hat{\phi}}}F_{D}\!\left(\bar{h}_{\star},-2n,\bar{h}_{\star}-\hat{h}_{\hat{\phi}},2(\hat{h}_{\hat{\phi}}+n),2\bar{h}_{\star};\frac{z-1}{z-x^{2}},\frac{z-1}{z},\frac{z-1}{z+x^{2}}\right)\,$, (31)
where $\bar{h}_{\star}=\frac{\Delta_{\star}+J_{\star}}{2}$ is the natural label for the restriction of the conformal group on the $P_{12}=0$ null line.
Note that, because of its second argument being a negative integer, the apparent three-variable Lauricella $F_{D}$ function considered above is actually a finite sum of two-variable Appell $F_{1}$ functions, which we define in Appendix B. For instance, if we consider a scalar leading-twist operator, we have $\bar{h}_{\star}=h_{\star}$ and $n=0$, which lead to
$\mathcal{G}(z,\bar{z},x)\stackrel{{\scriptstyle\bar{z}\to 1}}{{\simeq}}\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b_{\mathcal{O}_{\star}\hat{\phi}}(1-\bar{z})^{h_{\star}-2h_{\Phi}}(1-z)^{h_{\star}-2h_{\Phi}}z^{-h_{\star}+h_{\Phi}+\hat{h}_{\hat{\phi}}}\\ \times\left(\frac{1+x^{2}}{(x^{2}+z)(1+z)}\right)^{\hat{h}_{\hat{\phi}}}F_{1}\!\left(h_{\star},h_{\star}-\hat{h}_{\hat{\phi}},2\hat{h}_{\hat{\phi}},2h_{\star};\frac{z-1}{z},\frac{z-1}{z+x^{2}}\right)\,$, (32)
where the dependence on just two variables is manifest.
The defect-channel expansion
Having expressed the leading term in the bulk-channel expansion, our next task is to understand what objects we need to sum over in the defect channel side to reproduce this term. We use the Casimir singularity trick of Simmons-Duffin:2016wlq to simplify this task, leveraging the fact that the defect channel conformal blocks are eigenfunctions of Casimir operators of the parallel and transverse symmetry algebras. We have
$\frac{\mathcal{C}_{\parallel,i}^{2}\left(\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}\right)}{\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}}=(2\hat{h}_{i}+s)(2\hat{h}_{i}+s-p)\,,\qquad\frac{\mathcal{C}_{\perp,i}^{2}\left(\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}\right)}{\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}}=s(s+d-p-2)\,$, (33)
with $i=1,2$, where
$\mathcal{C}_{\parallel,i}^{2}=\frac{1}{2}\mathcal{T}_{AB}^{(i)}\mathcal{T}^{BA}_{(i)}\,,\qquad\mathcal{C}_{\perp,i}^{2}=\frac{1}{2}\mathcal{T}_{IJ}^{(i)}\mathcal{T}^{IJ}_{(i)}\,$, (34)
while $\mathcal{T}^{(i)}_{\mu\nu}=(P_{i})_{\mu}\partial_{P^{\nu}_{i}}-(P_{i})_{\nu}\partial_{P^{\mu}_{i}}$ is the differential action of the conformal generators on scalar fields.
We can now act with the Casimir operators on the full correlator in either the bulk channel or the defect channel expansions. From the action on the bulk channel (31) we can evince
$\frac{\mathcal{C}_{\parallel,i}^{2}\left(\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\mathcal{G}\right)}{\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\mathcal{G}}\propto(1-\bar{z})^{-2}\,,\qquad\frac{\mathcal{C}_{\perp,i}^{2}\left(\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\mathcal{G}\right)}{\Omega_{\Delta_{\Phi}\hat{\Delta}_{\hat{\phi}}}\mathcal{G}}\propto(1-\bar{z})^{-2}\,$, (35)
so the action of the Casimir operators on the correlator in the lightcone limit increases its divergent behavior for $\bar{z}\to 1$, while (33) remains valid block by block. Since this can be applied an arbitrary number of times [^6] Here, we are assuming generic values for the exponent of $(1-\bar{z})^{h_{\star}-2h_{\Phi}}$. Our arguments do not apply to the case where this exponent is a non-negative integer. making the correlator arbitrarily singular, this instructs us that the main contribution in the defect channel expansion has to come from blocks whose transverse spin scales as $s\propto(1-\bar{z})^{-1}$.
From Buric:2020zea , the general defect-channel conformal blocks are known exactly. These can be normalized in the OPE limit and expressed in our conventions as
[…]
where $\mathcal{N}=\frac{2^{-s}s!}{\left(\frac{q-2}{2}\right)_{s}}$ is a normalization factor, $C_{n}^{(m)}(x)$ is a Gegenbauer polynomial, the object
$F\!\left(v_{1},v_{2}\right)=F_{4}\!\left(\hat{h}_{1}\!+\!\hat{h}_{2}\!+\!s\!-\!\hat{h}_{\hat{\phi}},\,\hat{h}_{1}\!+\!\hat{h}_{2}\!+\!s\!-\!\hat{h}_{\hat{\phi}}\!+\!\tfrac{2-p}{2},\,2\hat{h}_{1}\!+\!s\!+\!\tfrac{2-p}{2},\,2\hat{h}_{2}\!+\!s\!+\!\tfrac{2-p}{2};v_{1},v_{2}\right)$ (37)
is an Appell $F_{4}$ function, and their set of cross-ratios is related to ours via
$\begin{split}v_{1}=-\frac{\left(1+x^{2}\right)^{2}z\bar{z}}{x^{2}\left(-1+z\bar{z}\right)^{2}}\,,\qquad v_{2}=-\frac{\left(z\bar{z}+x^{2}\right)^{2}}{x^{2}\left(-1+z\bar{z}\right)^{2}}\,,\qquad\cos\varphi=\frac{z+\bar{z}}{2(z\bar{z})^{\frac{1}{2}}}\,.\end{split}$ (38)
As we concluded from (35), in our analysis we just need the blocks at $\bar{z}\to 1$ with $s\propto(1-\bar{z})^{-1}$. We can retrieve these from (36) by rescaling $\bar{z}\to 1+\epsilon(\bar{z}-1)$ and $s\to\frac{s}{\epsilon}$, and sending $\epsilon\to 0$. The outcome of this procedure, detailed in Appendix D, is
$\hat{f}_{\hat{h}_{1},\hat{h}_{2},s}(z,\bar{z},x)\overunderset{s\rightarrow\infty}{\bar{z}\rightarrow 1}{\simeq}2^{-s}e^{-s(1-\bar{z})}x^{-2(\hat{h}_{\hat{\phi}}+\hat{h}_{1}-\hat{h}_{2})}(1+x^{2})^{\hat{h}_{\hat{\phi}}-2\hat{h}_{1}+2\hat{h}_{2}}z^{\hat{h}_{2}}(x^{2}+z)^{\hat{h}_{\hat{\phi}}+2\hat{h}_{1}-2\hat{h}_{2}}\\ \times(1-z)^{1-\frac{d}{2}}(1+z)^{-\hat{h}_{\hat{\phi}}}\,$. (39)
For a consistency check, we verified that this expression agrees (up to normalization) with the solution of the defect-channel Casimir equations with $\bar{z}\to 1+\epsilon(\bar{z}-1)$ and $s\to\frac{s}{\epsilon}$ in the limit $\epsilon\to 0$; in a similar fashion to the derivation of crossed-channel blocks in Kaviraj:2022wbw ; Harris:2024nmr .
The crossing equation
Combining the two expansions discussed above, we can recast the crossing equation at leading order in the $\bar{z}\to 1$ lightcone limit as
[…]
Since, as we already argued before, we are focusing on the contribution of large-spin operators, we can approximate the sum over transverse spins with an integral. This allows us to match the $\bar{z}$ dependence of the crossing equation by fixing the transverse-spin asymptotics of the OPE coefficients:
[…]
$\qquad b_{\Phi\widehat{\mathcal{O}}_{1}}b_{\Phi\widehat{\mathcal{O}}_{2}}\lambda_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}\hat{\phi}}\simeq C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}\frac{2^{s}s^{2h_{\Phi}-h_{\star}-1}}{\Gamma(2h_{\Phi}-h_{\star})}\,.$
We are then left with a simplified crossing equation
$\sum_{n}\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}(1-z)^{\bar{h}_{\star}-2h_{\Phi}}z^{-\bar{h}_{\star}+h_{\Phi}+\hat{h}_{\hat{\phi}}}(x^{2}-z)^{2n}(x^{2}+z)^{-2n-\hat{h}_{\hat{\phi}}}$ (42)
$\qquad\times\left(\frac{1+x^{2}}{1+z}\right)^{\hat{h}_{\hat{\phi}}}F_{D}\!\left(\bar{h}_{\star},-2n,\bar{h}_{\star}-\hat{h}_{\hat{\phi}},2(\hat{h}_{\hat{\phi}}+n),2\bar{h}_{\star};\frac{z-1}{z-x^{2}},\frac{z-1}{z},\frac{z-1}{z+x^{2}}\right)$
$=\sum_{\hat{h}_{1},\hat{h}_{2}}C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}\left(\frac{(1+x^{2})(x^{2}+z)}{x^{2}(1+z)}\right)^{\hat{h}_{\hat{\phi}}}\left(\frac{x^{2}+z}{x^{2}(1+x^{2})}\right)^{2\hat{h}_{1}-2\hat{h}_{2}}z^{\hat{h}_{2}}\left(1-z\right)^{1-\frac{d}{2}}\,$,
in which we are yet to constrain the sum over twists $\hat{h}_{1}$, $\hat{h}_{2}$ and to determine the form of the $C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}$ coefficients. Despite its simpler form, a complete matching of the two sides in (42) is an arduous task. For this reason, we will expand around additional simplifying limits in which this matching can be performed more easily. These limits correspond to the defect lightcone limit $z\to 0$, and the large separation limit $x\to\infty$, highlighted in Figure 6. In the next subsection, we will therefore focus on these two different simplifying limits to extract the coefficients for different infinite families of operators.
Figure 6: Here, we show the causal structure of the BBD setup in our fixed conformal frame. We highlight all the kinematic limits we consider in our bootstrap corresponding to the double lightcone limit $\bar{z}\to 1$ and $z\to 0$, together with the large separation limit $x\to\infty$.
3.2 Simplifying limit I: the defect lightcone OPE
As a first simplifying limit to tackle (42), we consider the $z\to 0$ limit that corresponds to sending one of the two bulk fields to be null separated from the defect, as highlighted in Figure 6. At leading order in this limit, the crossing equation (42) becomes
$\sum_{n}\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\,z^{h_{\Phi}}x^{-2\hat{h}_{\hat{\phi}}}F_{1}\!\left(\hat{h}_{\hat{\phi}},-2n,2(\hat{h}_{\hat{\phi}}+n),\bar{h}_{\star}+\hat{h}_{\hat{\phi}};\,x^{-2},-x^{-2}\right)\\ =\sum_{\hat{h}_{1},\hat{h}_{2}}C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}z^{\hat{h}_{2}}x^{2\hat{h}_{2}-2\hat{h}_{1}}\left(1+\frac{1}{x^{2}}\right)^{2\hat{h}_{2}-2\hat{h}_{1}}\,$, (43)
where $F_{1}$ is an Appell function of the first kind, as defined in (78). From the form of the $z\to 0$ crossing equation (43), it is straightforward to read off the allowed twists for the operators exchanged in the defect channel. On the one hand, matching the power of $z$ in both sides implies $\hat{h}_{2}=h_{\Phi}$. On the other hand, the $x$ dependence of the bulk channel can only be reproduced by the defect-channel twists $\hat{h}_{1}=h_{\Phi}+\hat{h}_{\hat{\phi}}+m$. We can then identify these contributions as coming from families of operators that approach the spectrum of generalized free-field theory (GFF) at large transverse spin:
$\hat{h}_{2}=h_{\Phi}$ $\Longrightarrow\quad\widehat{\mathcal{O}}_{2}=[\Phi]_{0,s}=\Box_{\perp}^{0}\partial^{i_{1}}_{\perp}\dots\partial^{i_{s}}_{\perp}\Phi\,$, (44)
$\hat{h}_{1}=h_{\Phi}+\hat{h}_{\hat{\phi}}+m$ $\Longrightarrow\quad\widehat{\mathcal{O}}_{1}=[\hat{\phi}\Phi]_{m,0,s}=\hat{\phi}\Box^{m}_{\parallel}\Box_{\perp}^{0}\partial^{i_{1}}_{\perp}\dots\partial^{i_{s}}_{\perp}\Phi+(\dots)\,$,
where we highlighted which derivatives are acting along parallel directions or transverse directions, and where the ellipsis stands for additional terms necessary to make the operators primaries. Note that, while these operators belong to the GFF spectrum, the double-twist operators $\hat{\phi}\Phi]_{m,0,s}$ do not couple directly to the bulk operator $\Phi$ in GFF: we have $b^{\text{GFF}}_{\Phi[\hat{\phi}\Phi]_{m,0,s}}=0$. This is also connected to the reason why no transverse boxes appear in (44). From the point of view of the defect, in fact, each transverse derivative operator $\Phi]_{m,s}$ defines an independent GFF sector that at leading order in spin does not interact with the other sectors. We can then interpret the terms we are probing in this limit as a leading-order perturbation away from GFF, where just the new bulk-to-defect coupling $b_{\Phi[\hat{\phi}\Phi]_{m,0,s}}$ is turned on.
To concretely match the two sides of the crossing equation, we can now expand around $x\to\infty$ and perform a matching at every order in $x^{-1}$. Including a rearrangement of the summation ranges, this leads to
$\sum_{n}\lambda_{\Phi\Phi\mathcal{O}_{\star}}\,b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\,\sum_{l=0}^{\infty}\frac{\bigl(\hat{h}_{\hat{\phi}}\bigr)_{l}\bigl(2\hat{h}_{\hat{\phi}}+2n\bigr)_{l}\,{}_{2}F_{1}\!\left[-l,-2n,-l-2n-2\hat{h}_{\hat{\phi}}+1;-1\right]}{l!\bigl(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}\bigr)_{l}}(-1)^{l}x^{-2l}$ (45)
$=\sum_{l=0}^{\infty}x^{-2l}\sum_{m=0}^{l}C_{[\hat{\phi}\Phi]_{m,0}[\Phi]_{0}}\,\binom{-2\hat{h}_{\hat{\phi}}-2m}{l-m}\,$,
which can now be recursively solved for the individual defect channel coefficients. This leads to the expression
$C_{[\hat{\phi}\Phi]_{m,0,s}[\Phi]_{0,s}}=\lambda_{\Phi\Phi\mathcal{O}_{\star}}\sum_{j=0}^{m}$ $\frac{\hat{h}_{\hat{\phi}}+j}{\hat{h}_{\hat{\phi}}+m}\binom{2\hat{h}_{\hat{\phi}}+2m}{m-j}\frac{(-1)^{j}\,\Gamma(2\bar{h}_{\star})\Gamma(\hat{h}_{\hat{\phi}}+j)}{j!\,\Gamma(\bar{h}_{\star})\Gamma(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}+j)}$ (46)
$\times\sum_{n=0}^{J/2}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\bigl(2\hat{h}_{\hat{\phi}}+2n\bigr)_{j}\,{}_{2}F_{1}(-j,-2n,-2\hat{h}_{\hat{\phi}}-j-2n+1;-1)\,$,
which via (41) determines the asymptotics of bulk-to-defect coefficients $b_{\Phi[\hat{\phi}\Phi]_{m,0,s}}$ at large transverse spin in terms of bulk CFT data and the free transverse-derivative couplings $b^{\text{GFF}}_{\Phi[\Phi]_{0,s}}$.
3.3 Simplifying limit II: large separation along the defect
To extract further data, we consider (42) in the $x\to\infty$ limit, namely when the field insertion in the defect $\hat{\phi}$ is at a large distance from the origin in the conformal frame of Figure 6.
The crossing equation at leading order becomes
$\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)(1-z)^{\bar{h}_{\star}-2h_{\Phi}}z^{-\bar{h}_{\star}+h_{\Phi}+\hat{h}_{\hat{\phi}}}\left(1+z\right)^{-\hat{h}_{\hat{\phi}}}{}_{2}F_{1}\!\left(\bar{h}_{\star},\bar{h}_{\star}-\hat{h}_{\hat{\phi}},2\bar{h}_{\star};\frac{z-1}{z}\right)$ (47)
$=\sum_{\hat{h}_{1},\hat{h}_{2}}C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}\left(x^{2}\right)^{\hat{h}_{2}-\hat{h}_{1}+\hat{h}_{\hat{\phi}}}z^{\hat{h}_{2}}\left(1-z\right)^{1-\frac{d}{2}}\left(1+z\right)^{-\hat{h}_{\hat{\phi}}}\,$,
where the sum over the tensor structure label $n$ in the bulk channel now only affects the bulk-to-defect coefficients $b_{\mathcal{O}_{\star}\hat{\phi}}^{(n)}$. It is straightforward to see that the independence of the LHS from $x$ implies twist additivity on the defect channel
$\hat{h}_{1}=\hat{h}_{2}+\hat{h}_{\hat{\phi}}\,$. (48)
To analyze this equation, we first simplify it by carrying some factors to the LHS
$\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\!\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)\frac{(1-z)^{\bar{h}_{\star}-2h_{\Phi}+\frac{d}{2}-1}}{z^{\bar{h}_{\star}-h_{\Phi}-\hat{h}_{\hat{\phi}}}}{}_{2}F_{1}\!\left(\bar{h}_{\star},\bar{h}_{\star}-\hat{h}_{\hat{\phi}},2\bar{h}_{\star};\frac{z-1}{z}\right)=\sum_{\hat{h}_{2}}C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}z^{\hat{h}_{2}}\,$, (49)
to then apply a hypergeometric identity valid for $\hat{h}_{\hat{\phi}}\notin\mathbb{Z}$, leading to
$\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\!\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)(1-z)^{\bar{h}_{\star}-2h_{\Phi}+\frac{d}{2}-1}z^{h_{\Phi}}\frac{\Gamma(2\bar{h}_{\star})}{\Gamma(\bar{h}_{\star})}\Biggl[\frac{\Gamma\left(\hat{h}_{\hat{\phi}}\right)}{\Gamma\left(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}\right)}{}_{2}F_{1}\!\left(\bar{h}_{\star},\bar{h}_{\star}-\hat{h}_{\hat{\phi}},1-\hat{h}_{\hat{\phi}};z\right)$ (50)
$+\frac{\Gamma\left(-\hat{h}_{\hat{\phi}}\right)}{\Gamma\left(\bar{h}_{\star}-\hat{h}_{\hat{\phi}}\right)}z^{\hat{h}_{\hat{\phi}}}{}_{2}F_{1}\!\left(\bar{h}_{\star},\bar{h}_{\star}+\hat{h}_{\hat{\phi}},1+\hat{h}_{\hat{\phi}};z\right)\Biggr]=\sum_{\hat{h}_{2}}C_{\widehat{\mathcal{O}}_{1}\widehat{\mathcal{O}}_{2}}z^{\hat{h}_{2}}\,.$
This equation can now be easily studied in an expansion around $z=0$. This immediately instructs us that the $x\to\infty$ limit involves contributions from two different families of operators in the defect channel, namely
[…]
and
$\mathrm{(II)}\quad\Longrightarrow\quad\begin{aligned} \hat{h}_{2}=h_{\Phi}+\hat{h}_{\hat{\phi}}+m\quad&\Longrightarrow\quad\widehat{\mathcal{O}}_{2}=[\hat{\phi}\Phi]_{0,m,s}=\hat{\phi}\Box^{0}_{\parallel}\Box_{\perp}^{m}\partial^{i_{1}}_{\perp}\dots\partial^{i_{s}}_{\perp}\Phi\,,\\ \hat{h}_{1}=h_{\Phi}+2\hat{h}_{\hat{\phi}}+m\quad&\Longrightarrow\quad\widehat{\mathcal{O}}_{1}=[\hat{\phi}\hat{\phi}\Phi]_{\vec{0},m,s}=\hat{\phi}^{2}\Box^{0}_{\parallel}\Box_{\perp}^{m}\partial^{i_{1}}_{\perp}\dots\partial^{i_{s}}_{\perp}\Phi\,.\end{aligned}$ (52)
The second family is a perhaps surprising find: it involves the “triple-twist” operators $\hat{\phi}\hat{\phi}\Phi]_{\vec{0},m,s}$, which can be thought of as double-twists of $\hat{\phi}\Phi]_{0,m,s}$ with another $\hat{\phi}$. The notation $\vec{0}$ is meant to stand for all the parallel quantum numbers that are not excited, including the degeneracy label that would normally be present for a parallel spin that is large enough.
We can now expand in $z$ explicitly the crossing equation and extract the data for the two families. For the first family, on the LHS we get
$\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\!\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)\frac{\Gamma(2\bar{h}_{\star})}{\Gamma(\bar{h}_{\star})}\frac{\Gamma\left(\hat{h}_{\hat{\phi}}\right)}{\Gamma\left(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}\right)}$ (53)
$\hskip 60.0pt\times\sum_{k,l}\frac{(-1)^{k}\left(\bar{h}_{\star}-2h_{\Phi}+\frac{d}{2}-k\right)_{k}\left(\bar{h}_{\star}\right)_{l}\left(\bar{h}_{\star}-\hat{h}_{\hat{\phi}}\right)_{l}}{k!l!\left(1-\hat{h}_{\hat{\phi}}\right)_{l}}z^{h_{\Phi}+k+l}$
$=\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\!\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)\frac{\Gamma(2\bar{h}_{\star})}{\Gamma(\bar{h}_{\star})}\frac{\Gamma\left(\hat{h}_{\hat{\phi}}\right)}{\Gamma\left(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}\right)}$
$\hskip 60.0pt\times\sum_{m}\sum_{k=0}^{m}\frac{(-1)^{k}\left(\bar{h}_{\star}-2h_{\Phi}+\frac{d}{2}-k\right)_{k}\left(\bar{h}_{\star}\right)_{m-k}\left(\bar{h}_{\star}-\hat{h}_{\hat{\phi}}\right)_{m-k}}{k!(m-k)!\left(1-\hat{h}_{\hat{\phi}}\right)_{m-k}}z^{h_{\Phi}+m}\,$,
which can be directly matched with the defect channel expansion to get
$C_{[\hat{\phi}\Phi]_{0,m,s}[\Phi]_{m,s}}=$ $\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\!\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)\frac{\Gamma(2\bar{h}_{\star})}{\Gamma(\bar{h}_{\star})}\frac{\Gamma\left(\hat{h}_{\hat{\phi}}\right)}{\Gamma\left(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}\right)}$ (54)
$\qquad\times\sum_{k=0}^{m}\frac{(-1)^{k}\left(\bar{h}_{\star}-2h_{\Phi}+\frac{d}{2}-k\right)_{k}\left(\bar{h}_{\star}\right)_{m-k}\left(\bar{h}_{\star}-\hat{h}_{\hat{\phi}}\right)_{m-k}}{k!(m-k)!\left(1-\hat{h}_{\hat{\phi}}\right)_{m-k}}\,.$
Similarly, the $z$-expansion for the second family allows us to determine
$C_{[\hat{\phi}\hat{\phi}\Phi]_{\vec{0},m,s}[\hat{\phi}\Phi]_{0,m,s}}=$ $\left(\lambda_{\Phi\Phi\mathcal{O}_{\star}}\!\sum_{n}b^{(n)}_{\mathcal{O}_{\star}\hat{\phi}}\right)\frac{\Gamma(2\bar{h}_{\star})}{\Gamma(\bar{h}_{\star})}\frac{\Gamma\left(-\hat{h}_{\hat{\phi}}\right)}{\Gamma\left(\bar{h}_{\star}-\hat{h}_{\hat{\phi}}\right)}$ (55)
$\quad\times\sum_{k=0}^{m}\frac{(-1)^{k}\left(\bar{h}_{\star}-2h_{\Phi}+\frac{d}{2}-k\right)_{k}\left(\bar{h}_{\star}\right)_{m-k}\left(\bar{h}_{\star}+\hat{h}_{\hat{\phi}}\right)_{m-k}}{k!(m-k)!\left(1+\hat{h}_{\hat{\phi}}\right)_{m-k}}\,.$
The approach we followed here can in principle be applied also to the subleading orders in $1/x$, for which the twist additivity condition (48) is modified by the addition of positive integers $m^{\prime}$ on the RHS. These can be interpreted as contributions coming from operators $\widehat{\mathcal{O}}_{1}$ with parallel boxes $\Box_{\parallel}^{m^{\prime}}$. Note that, due to the corrected twist additivity condition and the equal footing with which the two families of contributions $(\mathrm{I})$ and $(\mathrm{II})$ appear, it is natural to expect that the contributions of family $(\mathrm{II})$ will only pick up parallel boxes in the triple-twist operator. Due to the complicated nature of the bulk channel contribution, we did not work out closed form expressions at subleading orders in $1/x$ for the general case. This task can become easier in special cases such as $\mathcal{N}=4$ SYM where the leading-twist bulk contribution simplifies. We will discuss one such case in the next section.
4 Application: line defect in $\mathcal{N}=4$ SYM
We now present a concrete application of the bootstrap technology we developed above. Let us consider four-dimensional $\mathcal{N}=4$ SYM in the presence of a superconformal line defect, e.g. a $1/2$-BPS Maldacena-Wilson line  [^7] Actually, the results of this section do not depend on the dimension of the defect. We could in principle consider also a surface defect, but for simplicity we will assume the existence of a protected defect operator, the tilt operator, with dimension 1. There is no conceptual difficulty in considering surface defects.. We consider the $1/2$-BPS chiral primary operator
$\mathcal{O}_{20^{\prime}}(x,Y):=\mathrm{Tr}\left[\phi^{M}(x)\phi^{N}(x)Y_{M}Y_{N}\right]\,$, (56)
where $\phi^{M}$ are the six scalar fields in $\mathcal{N}=4$ SYM and $Y_{M}$ are complex null polarization vectors ($Y\cdot Y=0$) that project on the symmetric traceless representation ($\mathbf{20^{\prime}}$ or $[0,2,0]$) of the $\mathfrak{so}(6)$ $R$-symmetry algebra. This operator has protected scaling dimension $\Delta=2$ and is the superprimary of the stress tensor supermultiplet. Its two-point function in the presence of a line defect has been extensively studied in the literature Barrat:2020vch ; Barrat:2021yvp ; Gimenez-Grau:2023fcy . Here we would like to access a new class of CFT data by adding an operator on the defect. We choose the simplest possible defect scalar: the tilt operator.
The tilt operator is present in the spectrum of any conformal defect which breaks a global symmetry of the bulk CFT. In particular, if the bulk theory has a global symmetry with a current $J^{\mu}(x)$, the breaking of such symmetry is encoded in the equation
$\partial_{\mu}J^{\mu}(x)=\delta^{q}(x_{\perp})\hat{\phi}(x_{\parallel})$ (57)
where the operator $\hat{\phi}$ has protected scaling dimension $\hat{\Delta}=p$ (in particular $\hat{\Delta}=1$ for line defects). A line defect in $\mathcal{N}=4$ SYM breaks the $\mathfrak{psu}(2,2|4)$ superconformal algebra down to $\mathfrak{osp}(4^{*}|4)\supset\mathfrak{sl}(2,\mathbb{R})\oplus\mathfrak{so}(3)\oplus\mathfrak{so}(5)_{R}$. The preserved R-symmetry algebra is $\mathfrak{so}(5)\subset\mathfrak{so}(6)$ and we have five tilt operators, associated to the five broken generators, transforming in the vector representation of $\mathfrak{so}(5)$
$\hat{\phi}(x_{\parallel},y)=\phi^{m}y_{m}$ (58)
where $m=1,\dots,5$ and $y_{m}$ in this case is a $\mathfrak{so}(5)$ null polarization vector. This operator is also the superprimary of the displacement supermultiplet.
We are interested in the three-point function
$\expectationvalue{\mathcal{O}_{20^{\prime}}(x_{1},Y_{1})\mathcal{O}_{20^{\prime}}(x_{2},Y_{2})\hat{\phi}(x_{\parallel},y)}$ (59)
To carry out a complete analysis of this correlator one would need to study the constraints of supersymmetry finding superconformal Ward identities and superconformal blocks, a task which we leave for future work. Here we only exploit the knowledge of the bulk OPE selection rules to extract information about the large-spin defect spectrum.
The superselection rules for the fusion of two chiral primary operators are well known and, in the notation of Dolan:2002zh , they read
[…]
with $J=2j$. $\mathcal{B}^{\frac{1}{2},\frac{1}{2}}$ and $\mathcal{B}^{\frac{1}{4},\frac{1}{4}}$ represent $1/2$-BPS and $1/4$-BPS short supermultiplets respectively. The $\mathcal{C}$ supermultiplets are semishort, while the $\mathcal{A}$ labels long multiplets. For the case of the bulk two-point function analyzed in Barrat:2020vch ; Barrat:2021yvp ; Gimenez-Grau:2023fcy , one needs to select only the supermultiplets on the r.h.s. that have a non-vanishing one-point function in the presence of the conformal defect, i.e. the identity operator $\mathds{1}$, the $1/2$-BPS short multiplet $\mathcal{B}_{[0,2k,0]}$ with $k=1,2$, the semishort multiplets $\mathcal{C}_{[0,2,0],(j,j)}$ and long multiplets $\mathcal{A}_{[0,0,0],(j,j)}^{\Delta}$ with $\Delta\geq J+2$. In the lightcone bootstrap analysis, as we discussed earlier in this work, the leading contribution in the bulk channel comes from the lowest-twist exchanged operator, which in that case is just the identity operator.
For the BBD correlator we need to select all the supermultiplets that have a non-vanishing bulk-defect two-point function with the tilt operator inserted on the defect. A simple representation theory analysis shows that the allowed supermultiplets are
$\begin{split}\mathcal{B}_{[0,2,0](0,0)}^{\frac{1}{2},\frac{1}{2}}\times\mathcal{B}_{[0,2,0](0,0)}^{\frac{1}{2},\frac{1}{2}}\to&\,\mathcal{B}_{[0,2,0](0,0)}^{\frac{1}{2},\frac{1}{2}}+\mathcal{B}_{[0,4,0](0,0)}^{\frac{1}{2},\frac{1}{2}}+\sum_{l=0}^{\infty}\mathcal{C}_{[0,2,0](j,j)}^{\frac{1}{2},\frac{1}{2}}\\ &+\sum_{l=0}^{\infty}\mathcal{C}_{[1,0,1](j,j)}^{\frac{1}{4},\frac{1}{4}}+\sum_{\Delta,l}\mathcal{A}^{\Delta}_{[0,0,0](j,j)}\,.\end{split}$ (61)
We are interested in isolating the lowest-twist contribution in this OPE. Since the action of the supercharges cannot lower the twist of an operator, we only need to analyze the twist of the superprimary in each multiplet. The superprimary of the $1/2$-BPS multiplet $\mathcal{B}_{[0,2,0],(0,0)}^{\frac{1}{2},\frac{1}{2}}$, i.e. the stress-tensor multiplet, is the chiral primary $\mathcal{O}_{20^{\prime}}$ itself and has twist $2h=2$. The next short multiplet $\mathcal{B}_{[0,4,0],(0,0)}^{\frac{1}{2},\frac{1}{2}}$ contains a scalar superconformal primary whose conformal dimension is $\Delta=4$, therefore it is subleading in twist. Both semishort multiplets $\mathcal{C}_{[0,2,0](j,j)}^{\frac{1}{2},\frac{1}{2}}$ and $\mathcal{C}_{[1,0,1](j,j)}^{\frac{1}{4},\frac{1}{4}}$ contain a superconformal primary with conformal dimension $\Delta=4+2j$, i.e. subleading twist $2h=4$. Finally, for the long multiplet $\mathcal{A}^{\Delta}_{[0,0,0](j,j)}$ the unitarity bound imposes that $\Delta\geq 2+J$. Therefore, the only case we need to worry about is the threshold when the bound is saturated. In that case the long multiplet recombines in terms of shorter multiplets Dolan:2002zh
$\begin{split}&\mathcal{A}^{2}_{[0,0,0](0,0)}\simeq\mathcal{C}_{[0,0,0](0,0)}^{1,1}+\mathcal{B}_{[2,0,2](0,0)}^{\frac{1}{4},\frac{1}{4}}+\mathcal{D}_{[2,0,0](0,0)}^{\frac{1}{4},\frac{3}{4}}+\mathcal{D}_{[0,0,2](0,0)}^{\frac{3}{4},\frac{1}{4}}\,,\\ &\mathcal{A}^{2+2j}_{[0,0,0](j,j)}\simeq\mathcal{C}_{[0,0,0](j,j)}^{1,1}+\mathcal{C}_{[1,0,1](j-\frac{1}{2},j-\frac{1}{2})}^{\frac{1}{4},\frac{1}{4}}+\mathcal{C}_{[1,0,0](j-\frac{1}{2},j)}^{\frac{1}{4},\frac{3}{4}}+\mathcal{C}_{[0,0,1](j,j-\frac{1}{2})}^{\frac{3}{4},\frac{1}{4}}\,.\end{split}$ (62)
The superprimary is then contained in the $\mathcal{C}_{[0,0,0](j,j)}^{1,1}$ multiplet. These multiplets however are absent in an interacting theory because they contain higher-spin conserved currents Maldacena:2011jn . In other words, we do not expect long multiplets to hit the unitarity bound in an interacting theory. We can then conclude that the leading-twist contribution is given by the stress tensor supermultiplet.
The twist-two operators in the stress tensor multiplet are: the scalar $\mathcal{O}_{20^{\prime}}$ with $\Delta=2$, the R-symmetry current with $\Delta=3$ and $J=1$ and the stress-tensor with $\Delta=4$ and $J=2$. The current could give a non-vanishing contribution for specific parity-odd cases, but this is not the case for a line defect in four dimensions. In the following, we show explicitly how lightcone bootstrap techniques apply to the case of the dimension-two scalar, but the stress tensor case is completely analogous. Furthermore, we expect supersymmetric Ward identities to provide an explicit relation between $b_{\mathcal{O}\hat{\phi}}$ and $b_{T_{\mu\nu}\hat{\phi}}$, but we leave this analysis for the future.
We start from the crossing equation in the lightcone limit (42) and we fix conformal dimensions and spins for all the external operators and the lowest-twist bulk exchanged operator. In particular, we set $n=0$ (the only allowed tensor structure) and fix $\bar{h}_{\star}=h_{\Phi}=h_{\mathcal{O}_{20^{\prime}}}=1$, $\hat{h}_{\hat{\phi}}=\frac{1}{2}$, and $d=4$. The result is
[…]
where we defined $\lambda_{\mathcal{O}_{20^{\prime}}}\equiv\lambda_{\mathcal{O}_{20^{\prime}}\mathcal{O}_{20^{\prime}}\mathcal{O}_{20^{\prime}}}$. In the defect channel, from our discussion in Section 3.3, we expect two families of contributions: $(\mathrm{I})$ the exchange of a single twist operator with $\hat{h}_{2}=h_{\mathcal{O}_{20^{\prime}}}+m_{2}=1+m_{2}$ and a double twist with $\hat{h}_{1}=\hat{h}_{2}+\hat{h}_{\hat{\phi}}+m_{1}=\frac{3}{2}+m_{1}+m_{2}$, or $(\mathrm{II})$ the exchange of a double twist operator with $\hat{h}_{2}=h_{\mathcal{O}_{20^{\prime}}}+\hat{h}_{\hat{\phi}}+m_{2}=\frac{3}{2}+m_{2}$ and a triple twist888note that, for the case of a line defect, a number of odd parallel derivatives applied to the double-twist operators could produce objects with the same twist as the triple-twist operators considered here. We still keep the label in terms of triple-twist operators since for our results we never needed to specify the defect dimension $p$, so they would also apply to (subsectors of) surface defects contributions in $d=4$. with $\hat{h}_{1}=\hat{h}_{2}+\hat{h}_{\hat{\phi}}+m_{1}=2+m_{1}+m_{2}$. Inserting these values in the crossing equation we get
[…]
Let us start from the family $(\mathrm{I})$ of defect-channel contributions. The crossing equation becomes simply
[…]
By matching the expansions in $1/x$ and $z$ in this equation we find
$\begin{split}2\lambda_{\mathcal{O}_{20^{\prime}}}b_{\mathcal{O}_{20^{\prime}}\hat{\phi}}\frac{(-1)^{k}}{2k+1}=&\sum_{m_{1}=0}^{k}\sum_{m_{2}=\mathrm{max}(0,l-k+m_{1})}^{l}C_{[\hat{\phi}\mathcal{O}_{20^{\prime}}]_{m_{1},m_{2},s}[\mathcal{O}_{20^{\prime}}]_{m_{2},s}}\\ &\times\binom{2m_{1}+1}{l-m_{2}}\binom{-2m_{1}-1}{k-l+m_{2}-m_{1}}\end{split}$ (66)
from which one can recursively extract the coefficients $C_{[\hat{\phi}\mathcal{O}_{20^{\prime}}]_{m_{1},m_{2},s}[\mathcal{O}_{20^{\prime}}]_{m_{2},s}}$ for any value of $m_{1},m_{2}$. In particular, we find a very simple closed form for all the coefficients with $m_{1}\leq m_{2}$:
$C_{[\hat{\phi}\mathcal{O}_{20^{\prime}}]_{m_{1},m_{2},s}[\mathcal{O}_{20^{\prime}}]_{m_{2},s}}=2\lambda_{\mathcal{O}_{20^{\prime}}}b_{\mathcal{O}_{20^{\prime}}\hat{\phi}}\frac{(-1)^{m_{1}}}{2m_{1}+1}\,,\quad m_{1}\leq m_{2}\,.$ (67)
The procedure for the second family of operators is completely analogous. The crossing equation reduces to
[…]
and matching order by order the expansions in $1/x$ and $z$, we find a relation for the coefficients $C_{[\hat{\phi}\hat{\phi}\mathcal{O}_{20^{\prime}}]_{m_{1},m_{2},s}[\hat{\phi}\mathcal{O}_{20^{\prime}}]_{0,m_{2},s}}$ for any $m_{1},m_{2}$. In particular, we have that for any $j,l\in\mathbb{Z}_{>0}$ with $0\leq l\leq j$
[…]
while for $l<0$ we have
[…]
Solving these two equations allows one to extract all the coefficients. As before, there is actually a simple closed formula valid for any $m_{1}\leq m_{2}$:
$C_{[\hat{\phi}\hat{\phi}\mathcal{O}_{20^{\prime}}]_{m_{1},m_{2},s}[\hat{\phi}\mathcal{O}_{20^{\prime}}]_{0,m_{2},s}}=2\lambda_{\mathcal{O}_{20^{\prime}}}b_{\mathcal{O}_{20^{\prime}}\hat{\phi}}\frac{(-1)^{m_{1}+1}}{2m_{1}+1}\,,\quad m_{1}\leq m_{2}\,.$ (71)
5 Outlook
In this paper we focused on correlators involving two bulk operators and one defect operator, constraining the bulk-to-defect couplings of a bulk scalar $\Phi$ with both the double-twists $\Phi\hat{\phi}]_{m_{1},m_{2},s}$ or the triple-twists $\Phi\hat{\phi}\hat{\phi}]_{m_{1},m_{2},s}$ at large transverse spin. The main results for general CFTs are the closed-form expressions (41), (46), (54), and (54), while those specific to some correlators in $d=4$ theories such as $\mathcal{N}=4$ SYM are (67) and (71). The straightforward success of this approach leads to some natural directions for future exploration.
The appealing feature of the setup we considered here is the simplicity compared to higher-point correlators without defects. The insertion of operators on the defect affects the kinematics in a controlled way. For instance, the defect-channel conformal blocks preserve the factorized structure that was crucial in the derivation of the Lorentzian inversion formula Lemos:2017vnx and the dispersion relation Bianchi:2022ppi ; Barrat:2022psm . It is therefore natural to ask whether these techniques can be extended to our setup. This would give a rigorous and systematic way to understand the large-spin regime of the theory and it would provide a powerful computational technique for those theories that admit a perturbative expansion with some control over the spectrum of exchanged operators.
Throughout the paper we considered only defects with codimension greater than one. In some respect, the case of codimension-one defects (boundaries or interfaces) should be even simpler because the number of conformal cross ratios is further reduced. However, in this case there is no notion of transverse spin and the lightcone limit of the crossing equation should provide information on the defect spectrum at large conformal dimension. We leave a detailed analysis of this case for future work.
Finally, another natural direction is the extension of lightcone bootstrap techniques to other higher-point correlators. We skecthed some possibilities in Figure 1. In particular, it would be very interesting to analyze the BBDD and the BBB case, which would give access to a whole new class of operators. Given that the triple-twist class of contributions we observed in the BBD case was somewhat unexpected in this setting, it would be indeed interesting to address which kind of new contributions one has access to when increasing the number of field insertions. Moreover, one may wonder whether the simplicity we observe in the structure of the defect conformal blocks survives after additional insertions, and whether some extrapolations to a large numbers of insertions can be made. We plan to address this in the future.
Acknowledgements.
The authors would like to thank M. Lemos, M. Meineri and V. Schomerus for useful discussions. LB’s and LQ’s research is partially supported by the MUR PRIN contract 2022N9CTAE "Constraining strongly coupled quantum field theories using symmetry". LB and AM would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme "Quantum Field Theory with Boundaries, Impurities, and Defects" where work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1. The authors participate in the project HORIZON-MSCA-2023-SE-01-101182937HeI.
Appendix A Embedding space formalism
Let us consider a $d$-dimensional Euclidean CFT living in $\mathbb{R}^{d}$, with a $p$-dimensional flat conformal defect. We work in the embedding space formalism, which has the advantage that conformal transformations act linearly on the coordinates. In particular, the embedding space realizes the $d$-dimensional flat space as a subset of the projective null cone in $\mathbb{R}^{d+1,1}$. Indices in embedding space are $A,B,\ldots=(+,-,1,\ldots,p)$ for the $SO(p+1,1)$ defect conformal group and $I,J,\ldots=(1,\ldots,d-p)$ for the transverse rotation part SO(dp)SO(d-p). We will also use the definition q=dpq=d-p to indicate the codimension. The defect location is such that all components of the transverse indices $I,J,\ldots$ are zero. The two indices are associated with two scalar products
$P\!\bullet\!Q=P^{A}\eta_{AB}Q^{B}\,,\qquad P\!\circ\!Q=P^{I}\delta_{IJ}Q^{J}\,.$ (72)
For spinning operators, we will use the index-free notation. In particular, a symmetric traceless bulk operator $\mathcal{O}_{\Delta,J}(P,Z)$ is obtained by contracting the indices with a polarization vector $Z$ such that $P^{2}=0$, $Z^{2}=0$, $Z\cdot P=0$. Splitting in parallel and orthogonal directions we have
$P\!\bullet\!P=-P\!\circ\!P\,,\qquad Z\!\bullet\!Z=-Z\!\circ\!Z\,,\qquad Z\!\bullet\!P=-Z\!\circ\!P\,.$ (73)
As usual, due to the transversality condition (gauge invariance), $Z$ can only appear in correlation functions through a wedge product. In this case, though, only some mixed components are independent
$C^{AI}=(P\wedge Z)^{AI}=P^{A}Z^{I}-P^{I}Z^{A}\,.$ (74)
For example, the structure of a one-point function of a primary bulk operator $\Phi_{\Delta,J}(P,Z)$ can be constructed with this tensor structure $C^{AI}$. In particular, the $Z$ polarization can only appear from $C_{AI}C^{AI}$, which implies that only even spin operators acquire a non-vanishing one-point function (at least for parity-preserving theories). Fixing homogeneities appropriately, we have
$\expectationvalue{O_{\Delta,J}(P,Z)}=a_{\mathcal{O}}\frac{\left(C_{AI}C^{AI}\right)^{\frac{J}{2}}}{\left(P\!\circ\!P\right)^{\Delta+J}}\,$, (75)
where the normalization $a_{\mathcal{O}}$ is part of the defect CFT data.
On the other hand, defect operators are represented with a hat $\widehat{\mathcal{O}}_{\Delta,j,s}(P,Z,W)$ and characterized by three embedding space vectors $P^{A}$, $Z^{A}$, $W^{I}$ associated to spacetime, parallel spin and orthogonal spin respectively. They are constrained by
$P\!\bullet\!P=0\,,\qquad Z\!\bullet\!Z=0\,,\qquad Z\!\bullet\!P=0\,,\qquad W\!\circ\!W=0\,.$ (76)
Appendix B Generalized hypergeometric functions
Throughout the paper, we encounter several special functions, we report some definitions here for the reader’s convenience.
The Lauricella $F_{D}$ function, a generalized multivariable hypergeometric function, can be defined as
[…]
$=\sum_{(n_{1},n_{2},n_{3})\in\mathbb{N}_{0}^{3}}\!\frac{(a)_{n_{1}+n_{2}+n_{3}}\prod_{i=1}^{3}(b_{i})_{n_{i}}}{(c)_{n_{1}+n_{2}+n_{3}}\prod_{i=1}^{3}n_{i}!}x_{1}^{n_{1}}x_{2}^{n_{2}}x_{3}^{n_{3}}\,.$
Note that, similarly to what happens in the case of the standard hypergeometric functions, if one of the $b_{i}$ is a negative integer, the sum over the associated $n_{i}$ truncates to a finite number of terms. This is precisely the case for the Lauricella $F_{D}$ that appears in our bulk-channel blocks in Section 3. In this case, the function becomes a finite sum of Appell $F_{1}$ functions, a class of hypergeometric functions with two variables that we define below.
The Appell function $F_{1}$ is defined in its series and integral representation as
$\begin{split}F_{1}(a,b_{1},b_{2},c,z_{1},z_{2})&=\sum_{m,n=0}^{\infty}\frac{(a)_{m+n}(b_{1})_{m}(b_{2})_{n}}{(c)_{m+n}\,m!\,n!}z_{1}^{m}z_{2}^{n}\\ &=\frac{\Gamma(c)}{\Gamma(c-a)\,\Gamma(a)}\int_{0}^{1}\mathrm{d}t\,\frac{t^{a-1}(1-t)^{c-a-1}}{(1-t\,z_{1})^{b_{1}}(1-t\,z_{2})^{b_{2}}}\,.\end{split}$ (78)
This function is both relevant in the decomposition of the Lauricella function we encounter in the bulk channel, and in the first simplifying limit we consider in Section 3.2.
In the case of defect-channel conformal blocks, the general solution is expressed in terms of an Appell function $F_{4}$, which we can define as
$\begin{split}F_{4}(a,&b,c_{1},c_{2},z_{1},z_{2})=\sum_{m,n=0}^{\infty}\frac{(a)_{m+n}(b)_{m}(b_{2})_{m+n}}{(c_{1})_{m}(c_{2})_{n}\,m!\,n!}z_{1}^{m}z_{2}^{n}\\ &=\frac{\Gamma(c_{1})\Gamma(c_{2})}{\Gamma(a)\Gamma(b)\Gamma(c_{1}-a)\Gamma(c_{2}-b)}\\ &\times\int_{0}^{1}\int_{0}^{1}\mathrm{d}u\,\mathrm{d}v\,\frac{u^{a-1}v^{b-1}(1-u)^{c_{1}-a-1}(1-v)^{c_{2}-b-1}}{(1-z_{1}\,u)^{c_{1}+c_{2}-a-1}(1-z_{2}\,v)^{c_{1}+c_{2}-b-1}(1-z_{1}\,u-z_{2}\,v)^{a+b-c_{1}-c_{2}+1}}\,.\end{split}$ (79)
Note that the integral representation is now a non-factorized double integral.
It is also useful to recall the following reduction formula for the $F_{4}$:
$\begin{split}F_{4}&\left(a,b,c_{1},c_{2},x_{1}(1-x_{2}),x_{2}(1-x_{1})\right)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}(a+b-c_{1}-c_{2}+1)_{k}}{(c_{1})_{k}(c_{2})_{k}k!}x_{1}^{k}x_{2}^{k}\,\\ &\times{}_{2}F_{1}(a+k,b+k,c_{1}+k,x_{1}){}_{2}F_{1}(a+k,b+k,c_{2}+k,x_{2})\,.\end{split}$ (80)
Appendix C Bulk-channel conformal blocks
C.1 Bulk-bulk two-point function
In this appendix, we review the bulk conformal blocks for the case of the bulk two-point function following the argument in Billo:2016cpy . In the bulk channel (8),
[…]
there is a sum over bulk primary operators appearing in the $\Phi\times\Phi$ OPE with a non-vanishing one-point function. As we explained in appendix A, we can restrict the sum to even J.
[…]
$\mathcal{T}_{MN}^{(i)}=P^{(i)}_{M}\frac{\partial}{\partial P^{N}_{(i)}}-P_{N}^{(i)}\frac{\partial}{\partial P^{M}_{(i)}}\,$, (82)
where $M,N$ are indices in the embedding space $M,N=(+,-,1,\ldots,d)$ for the full conformal group $SO(d+1,1)$ and with $i=1,2$ corresponding to the two bulk insertions. Then, the Casimir differential equation becomes
[…]
$f_{\Delta,J}(\xi,\varphi)=\xi^{\frac{\Delta-J}{2}}\left(\sin^{J}\varphi\,\,{}_{2}F_{1}\left(\frac{\Delta+J}{4},\frac{\Delta+J}{4},\frac{\Delta+J+1}{2},\sin^{2}\varphi\right)+O(\xi)\right)\,$, (84)
where we see that the lightcone limit organizes in terms of operators with lowest twist $\tau=\Delta-J$. Let us express this in terms of the cross ratios $z,\bar{z}$ defined in (10) as done in Lemos:2017vnx . The lightcone limit corresponds now to taking the limit $\bar{z}\to 1$ and we can express the bulk conformal block as in (17) with
$\tilde{f}_{\Delta+J}(z)=2^{-J}(1-z)^{\frac{\Delta+J}{2}}\,{}_{2}F_{1}\left(\frac{\Delta+J}{2},\frac{\Delta+J}{2},\Delta+J,1-z\right)\,.$ (85)
C.2 Bulk-bulk-defect three-point function
Let us now consider the bulk channel for the bulk-bulk-defect setup. In the bulk channel, we can use the $P_{1}\cdot P_{2}\to 0$ lightcone OPE Ferrara:1971zy ; Ferrara:1971tq ; Ferrara:1972cq to decompose the correlator in terms of bulk-to-defect correlators. Using the notation of (Kaviraj:2022wbw, , Appendix A) up to normalization of embedding space vectors, the scalar-scalar lightcone OPE can be written as
[…]
with $\tilde{P}=P_{2}-t(P_{2}-P_{1})$, $\tilde{Z}=P_{2}-P_{1}$. This equation is compatible with the normalization of two- and three-point functions.
Using the lightcone OPE (86), we can work out an expression for the lightcone bulk conformal blocks for the Bulk-Bulk-Defect correlator we are interested in. Inserting this lightcone OPE formula in our correlation function, we obtain an integral of a bulk-to-defect two-point function
$\expectationvalue{\!\mathcal{O}_{\Delta,J}(\tilde{P},\tilde{Z})\hat{\phi}(P_{3})\!}=\frac{\left(4\tilde{C}_{1}^{AI}P^{3}_{A}\tilde{P}_{I}\right)^{J}}{\left(-2\tilde{P}\!\bullet\!P_{3}\right)^{\hat{\Delta}_{\hat{\phi}}+J}\!\left(\tilde{P}\!\circ\!\tilde{P}\right)^{\frac{\Delta-\hat{\Delta}_{\hat{\phi}}+J}{2}}}\sum_{n}b_{\mathcal{O}\hat{\phi}}^{(n)}\!\left(\!\frac{\tilde{C}^{AI}\tilde{C}_{AI}\left(\tilde{P}\!\bullet\!P_{3}\right)^{2}}{2\left(\tilde{C}^{AI}P^{3}_{A}\tilde{P}_{I}\right)^{2}}\!\right)^{\frac{J}{2}-n}\hskip-5.0pt\,$, (87)
where $n=0,1,\dots,J/2$ and $J\in 2\mathbb{N}$. With $\tilde{P}_{1}=P_{2}-t(P_{2}-P_{1})$ and $\tilde{Z}_{1}=P_{2}-P_{1}$ this becomes
$\begin{split}\sum_{n}&b^{(n)}_{\mathcal{O}\hat{\phi}}\frac{\left[4\left(P_{1}\!\bullet\!P_{3}\right)\left(P_{2}\!\circ\!P_{2}\right)-4\left(P_{2}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{2}\right)\right]^{2n}}{\left(-2P_{2}\!\bullet\!P_{3}\right)^{\hat{\Delta}_{\hat{\phi}}+2n}\left(P_{2}\!\circ\!P_{2}\right)^{\frac{\Delta-\hat{\Delta}_{\hat{\phi}}+J}{2}}}\left[\left(P_{1}\!\circ\!P_{2}\right)^{2}-\left(P_{1}\!\circ\!P_{1}\right)\left(P_{2}\!\circ\!P_{2}\right)\right]^{\frac{J}{2}-n}\\ &\times\left[1-t\left(1-\frac{\left(P_{2}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{1}\right)-\left(P_{1}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{2}\right)}{\left(P_{2}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{2}\right)-\left(P_{1}\!\bullet\!P_{3}\right)\left(P_{2}\!\circ\!P_{2}\right)}\right)\right]^{2n}\left[1-t\left(1-\frac{P_{1}\!\bullet\!P_{3}}{P_{2}\!\bullet\!P_{3}}\right)\right]^{-\hat{\Delta}_{\hat{\phi}}-2n}\\ &\times\left[1-t\frac{\left(P_{2}\!\circ\!P_{2}\right)-\left(P_{1}\!\circ\!P_{2}\right)-\sqrt{\left(P_{1}\!\circ\!P_{2}\right)^{2}-\left(P_{1}\!\circ\!P_{1}\right)\left(P_{2}\!\circ\!P_{2}\right)}}{P_{2}\!\circ\!P_{2}}\right]^{\frac{\hat{\Delta}_{\hat{\phi}}-\Delta-J}{2}}\\ &\times\left[1-t\frac{\left(P_{2}\!\circ\!P_{2}\right)-\left(P_{1}\!\circ\!P_{2}\right)+\sqrt{\left(P_{1}\!\circ\!P_{2}\right)^{2}-\left(P_{1}\!\circ\!P_{1}\right)\left(P_{2}\!\circ\!P_{2}\right)}}{P_{2}\!\circ\!P_{2}}\right]^{\frac{\hat{\Delta}_{\hat{\phi}}-\Delta-J}{2}}\,.\end{split}$ (88)
The square root can be evaluated as
$\sqrt{\left(P_{1}\!\circ\!P_{2}\right)^{2}-\left(P_{1}\!\circ\!P_{1}\right)\left(P_{2}\!\circ\!P_{2}\right)}=\sqrt{\left[\left(P_{1}-P_{2}\right)\!\circ\!P_{2}\right]^{2}}=\left(P_{1}-P_{2}\right)\!\circ\!P_{2}\,$, (89)
valid as long as $(P_{1}-P_{2})\!\circ\!(P_{1}-P_{2})=0$ and where the sign of the square root is chosen compatibly with the standard conformal frame we consider. Then, the bulk-defect two-point function becomes
$\begin{split}\sum_{n}b^{(n)}_{\mathcal{O}\hat{\phi}}&\frac{\left[4\left(P_{1}\!\bullet\!P_{3}\right)\left(P_{2}\!\circ\!P_{2}\right)-4\left(P_{2}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{2}\right)\right]^{2n}}{\left(-2P_{2}\!\bullet\!P_{3}\right)^{\hat{\Delta}_{\hat{\phi}}+J-2n}\left(P_{2}\!\circ\!P_{2}\right)^{\frac{\Delta+J-\hat{\Delta}_{\hat{\phi}}}{2}}}\left[\left(P_{1}\!\circ\!P_{2}\right)^{2}-\left(P_{1}\!\circ\!P_{1}\right)\left(P_{2}\!\circ\!P_{2}\right)\right]^{\frac{J}{2}-n}\\ &\times\left[1-t\left(1-\frac{\left(P_{2}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{1}\right)-\left(P_{1}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{2}\right)}{\left(P_{2}\!\bullet\!P_{3}\right)\left(P_{1}\!\circ\!P_{2}\right)-\left(P_{1}\!\bullet\!P_{3}\right)\left(P_{2}\!\circ\!P_{2}\right)}\right)\right]^{2n}\\ &\times\left[1-t\left(1-\frac{P_{1}\!\bullet\!P_{3}}{P_{2}\!\bullet\!P_{3}}\right)\right]^{-\hat{\Delta}_{\hat{\phi}}-2n}\left[1-t\left(2-2\frac{P_{1}\!\circ\!P_{2}}{P_{2}\!\circ\!P_{2}}\right)\right]^{\frac{\hat{\Delta}_{\hat{\phi}}-\Delta-J}{2}}.\end{split}$ (90)
If we insert this in the lightcone integral, we get the integral representation for a Lauricella $F_{D}$ function as defined in (77). This allows us to express the correlator expanded in the bulk channel as
[…]
where
[…]
with $h=\frac{\Delta-J}{2}$ being the half twist of the bulk exchanged operators. Let us also define for simplicity $\bar{h}=h+J=\frac{\Delta+J}{2}$ and $\hat{h}_{\hat{\phi}}=\frac{\Delta_{\hat{\phi}}}{2}$. We notice that one of the sums of the Lauricella truncates, due to 2n-2n in the second argument being a negative integer. Fixing the positions of fields as in the conformal frame depicted in Figure 5, the bulk-channel expansion of the correlator (91) becomes:
$\langle\Phi(P_{1})\Phi(P_{2})\hat{\phi}(P_{3})\rangle\stackrel{{\scriptstyle\text{frame}}}{{=}}\sum_{\mathcal{O},n}\lambda_{\Phi\Phi\mathcal{O}}\,b^{(n)}_{\mathcal{O}\hat{\phi}}(1-\bar{z})^{h-2h_{\Phi}}(1-z)^{\bar{h}-2h_{\Phi}}z^{-\bar{h}+\hat{h}_{\hat{\phi}}}\frac{(x^{2}-z)^{2n}}{(x^{2}+z)^{2n+2\hat{h}_{\hat{\phi}}}}\\ \times F_{D}\!\left(\bar{h},-2n,\bar{h}-\hat{h}_{\hat{\phi}},2(\hat{h}_{\hat{\phi}}+n),2\bar{h};\frac{z-1}{z-x^{2}},\frac{z-1}{z},\frac{z-1}{z+x^{2}}\right)\,.$ (93)
Note that this expression, which we derived from (91), is not manifestly symmetric under $1\leftrightarrow 2$. The lightcone OPE is, however, a symmetric operation, so there must exist Lauricella identities that allow us to express the conformal blocks with equivalent representations. One such representation is obtained by exchanging $P_{1}\leftrightarrow P_{2}$ in (91), obtaining
$\langle\Phi(P_{1})\Phi(P_{2})\hat{\phi}(P_{3})\rangle\stackrel{{\scriptstyle\text{frame}}}{{=}}\sum_{\mathcal{O},n}\lambda_{\Phi\Phi\mathcal{O}}\,b^{(n)}_{\mathcal{O}\hat{\phi}}\left(1-\bar{z}\right)^{h-2h_{\phi}}\frac{(1-z)^{\bar{h}-2h_{\phi}}\left(1-x^{2}\right)^{2n}}{\left(x^{2}+1\right)^{2\hat{h}_{\hat{\phi}}+2n}}\\ \times F_{D}\!\left(\bar{h},-2n,\bar{h}-\hat{h}_{\hat{\phi}},2(\hat{h}_{\hat{\phi}}+n),2\bar{h},\frac{z-1}{x^{2}-1},1-z,\frac{1-z}{x^{2}+1}\right)\,.$ (94)
These entail non-trivial identities between Lauricella functions, which allow for analytic continuations in regimes where one of the representations is problematic.
As a further consistency check for our bulk conformal block, let us consider the case in which the external scalar defect operator $\hat{\phi}(P_{3})$ is just the identity. The expectation is to reproduce the bulk conformal block for the bulk two-point function in the lightcone limit (85). To impose the external defect operator to be the identity we need to fix $\hat{h}_{\hat{\phi}}=0$ and trivialize the tensor structure label to $n=0$. The bulk channel in (93) then loses all dependence on the $x$ cross-ratio and reduces to
[…]
where $b_{\mathcal{O}\hat{\mathds{1}}}=a_{\mathcal{O}}$ is just the one-point function coefficient. Using the following identity for the hypergeometric function $_{2}F_{1}$
$_{2}F_{1}\left(a,c-b,c,\frac{z}{z-1}\right)=(1-z)^{a}\,{}_{2}F_{1}(a,b,c,z),\quad z\notin(1,\infty),\,$ (96)
then, we can reduce the bulk channel as
[…]
where on the second line we multiplied and divided for $(z\bar{z})^{-\Delta_{\Phi}/2}$. It is straightforward to notice that this expression, with $\tilde{f}_{\Delta+J}(z)$ defined in (85), is equivalent [^9] Actually, our result differs with the one in Lemos:2017vnx for a factor $2^{-J}$, which we absorb it in the definition of the bulk-to-defect OPE coefficient $b_{\mathcal{O}\hat{\phi}}^{(n)}$. to the bulk channel for the bulk two-point function case (8), (14), (17) in the lightcone limit.
Appendix D Large spin expansion of defect-channel blocks
In this appendix, we will show how to perform the large transverse spin expansion of the defect conformal blocks for the cases of the bulk-bulk two-point function and the bulk-bulk-defect three-point function.
Bulk-bulk blocks in the defect channel
Let us first consider the bulk-bulk conformal block in the defect channel obtained analytically in Billo:2016cpy , and let us write it in the convention of Lemos:2017vnx :
$\hat{f}_{\hat{h},s}(z,\bar{z})=z^{\hat{h}}\bar{z}^{\hat{h}+s}{}_{2}F_{1}\!\left(-s,\frac{q}{2}-1,2-\frac{q}{2}-s;\frac{z}{\bar{z}}\right){}_{2}F_{1}\!\left(2\widehat{h}+s,\frac{p}{2},2\widehat{h}+s+1-\frac{p}{2};z\bar{z}\right)\,.$ (98)
In the lightcone bootstrap, one is interested in reproducing the low-twist bulk exchange in the lightcone limit $\bar{z}\rightarrow 1$ in the defect channel. For the simpler case of a bulk two-point function, the lowest-twist contribution in the bulk OPE is the identity. Here, the crossing equation takes the form
$1=\lim_{\bar{z}\rightarrow 1}\left(\frac{(1-z)(1-\bar{z})}{\sqrt{z\bar{z}}}\right)^{\Delta_{\Phi}}\sum_{\hat{h},s}b_{\hat{h},s}^{2}\hat{f}_{\hat{h},s}(z,\bar{z})\,.$ (99)
Since each conformal block is analytic in the region of interest $1-\bar{z}\ll z<1$, then in the lightcone limit $(1-\bar{z})\rightarrow 0$ one finds that $(1-\bar{z})^{\Delta_{\Phi}}\hat{f}_{\hat{h},s}=0$ for fixed $\hat{h}$ and $s$, so the sum on the RHS does not converge uniformly. As shown in previous works Fitzpatrick:2014vua ; Simmons-Duffin:2016wlq ; Komargodski:2012ek ; Lemos:2017vnx , the region of the defect spectrum responsible for reproducing the pole in $1-\bar{z}$ is that of fixed transverse twist $2\hat{h}$ and large transverse spin $s$. A subtlety is that the large transverse spin $s$ limit and the lightcone $1-\bar{z}$ limit cannot be taken independently. Indeed, there is an interplay between these two limits which highlights the spin range responsible for reproducing the bulk OPE limit Lemos:2017vnx
$s\sim\frac{1}{1-\bar{z}}\,$, (100)
which should be contrasted with the usual quadratic relation $l^{2}\sim 1/(1-\bar{z})$ that appears in the homogenous four-point function (here $l$ denotes the Lorentz spin in $SO(d)$). This also holds in the bulk-bulk-defect three-point function that is the focus of this paper. Indeed, another way to see this behavior is through the Casimir-singularity trick Simmons-Duffin:2016wlq as shown in Section 3.
To perform the matching in the crossing equation (99), one therefore needs the large transverse spin expansion of the defect conformal block (98). Since the transverse spin appears in two distinguished hypergeometric functions, it is useful to isolate the different $s$-dependent pieces. Let us first consider the second hypergeometric function with argument $z\bar{z}$. For this function, we can apply a well-known identity of the $_{2}F_{1}$ hypergeometric functions [^10] Another convenient way to perform an expansion of $_{2}F_{1}$ at large parameters is to use its integral representation and apply a saddle-point approximation. See, for example, Appendix A in Bissi:2024wur . analogous to the one used in (96)
$_{2}F_{1}\left(c-a,b,c,\frac{z}{z-1}\right)=(1-z)^{b}\,{}_{2}F_{1}(a,b,c,z)\,,\quad z\notin(1,\infty),\,$ (101)
which allows us to shift the large-parameter $s$ dependence just in the third argument of the hypergeometric function. At this point, the hypergeometric function collapses to just a single term due to the large $s$ limit:
$\begin{split}{}_{2}F_{1}\!\left(2\widehat{h}+s,\frac{p}{2},2\widehat{h}+s+1-\frac{p}{2};z\bar{z}\right)&=(1-z\bar{z})^{-\frac{p}{2}}{}_{2}F_{1}\!\left(1-\frac{p}{2},\frac{p}{2},2\widehat{h}+s+1-\frac{p}{2};\frac{z\bar{z}}{z\bar{z}-1}\right)\\ &\overunderset{s\to\infty}{\bar{z}\to 1}{\sim}(1-z)^{-\frac{p}{2}}\,.\end{split}$ (102)
[…]
$s\rightarrow\frac{s}{\epsilon}\,,\quad\bar{z}\to 1+\epsilon(\bar{z}-1)\,,\quad\epsilon\rightarrow 0\,.$ (103)
With this prescription, the hypergeometric function we are considering has the simple leading behavior
$\lim_{\epsilon\rightarrow 0}{}_{2}F_{1}\left(-\frac{s}{\epsilon},\frac{q}{2}-1,2-\frac{q}{2}-\frac{s}{\epsilon},\frac{z}{1+\epsilon(1-\bar{z})}\right)\sim(1-z)^{1-q/2}\,.$ (104)
Finally, in the lightcone limit $\bar{z}\to 1$ and at large transverse spin, the defect-channel conformal block in (98) becomes
$\hat{f}_{\hat{h},s}(z,\bar{z})\overunderset{s\to\infty}{\bar{z}\to 1}{\sim}e^{-s(1-\bar{z})}z^{\hat{h}}(1-z)^{1-\frac{d}{2}}\,.$ (105)
Bulk-bulk-defect blocks in the defect channel
Let us now consider the case of the bulk-bulk-defect three-point function. We want to study the large transverse spin behavior of the defect conformal blocks (36). Since the conformal blocks factorize, it is once again more convenient to expand each term independently. Moreover, since the transverse part of the defect-channel blocks matches that of the bulk-bulk case, we can simply re-use (104) to extract its asymptotics.
We are left with the analysis of the asymptotic behavior of the Appell function $F_{4}$ (37) that appears in the defect-channel block. In the lightcone limit $1-\bar{z}\rightarrow 0$ the two variables $(v_{1},v_{2})$ remain finite, while all four parameters of $F_{4}$ scale with $s$ in the same way. For the large-$s$ expansion, it is convenient to use the reduction formula introduced in (80). This representation expresses the Appell function as an infinite sum of products of two Gauss hypergeometric functions, and it is particularly useful for performing large-parameter asymptotic expansions. In terms of the cross-ratios $(x,z,\bar{z})$, the variables $x_{1},x_{2}$ read
$x_{1}=\frac{(1+x^{2})z\bar{z}}{x^{2}(-1+z\bar{z})}\,,\quad x_{2}=\frac{x^{2}+z\bar{z}}{-1+z\bar{z}}\,.$ (106)
Since all Appell parameters scale in the same way at large transverse spin, it is convenient to separate their finite contributions:
$a=a^{(0)}+s\,,\quad b=b^{(0)}+s\,,\quad c_{1}=c_{1}^{(0)}+s\,,\quad\quad c_{2}=c_{2}^{(0)}+s\,.$ (107)
It is straightforward to see that the prefactor formed by the ratios of Pochhammer symbols is finite in the $s\rightarrow\infty$ limit
$\frac{(a^{(0)}+s)_{k}(b^{(0)}+s)_{k}(a^{(0)}+b^{(0)}-c_{1}^{(0)}-c_{2}^{(0)}+1)_{k}}{(c_{1}^{(0)}+s)_{k}(c_{2}^{(0)}+s)_{k}k!}\overset{s\to\infty}{\sim}\frac{(a^{(0)}+b^{(0)}-c_{1}^{(0)}-c_{2}^{(0)}+1)_{k}}{k!}\,.$ (108)
The two Gauss hypergeometric functions appearing on the RHS of (80) have all their parameters large. Applying the following hypergeometric identity
$_{2}F_{1}(a,b,c;z)=(1-z)^{c-a-b}{}_{2}F_{1}(c-a,c-b,c,z)\,$, (109)
we shift the large-parameter dependence into the prefactor $(1-z)^{c-a-b}$. In our case, the large parameter appears in the third argument of the hypergeometric function on the RHS, hence the hypergeometric function trivializes. Combining this observation with the large-$s$ behavior of the Pochhammer prefactor in (108), we obtain the following leading asymptotic form of the Appell function $F_{4}$ in (80):
$\begin{split}&\lim_{s\rightarrow\infty}F_{4}\left(a^{(0)}+s,b^{(0)}+s,c_{1}^{(0)}+s,c_{2}^{(0)}+s,x_{1}(1-x_{2}),x_{2}(1-x_{1})\right)\\ &\sim\sum_{k=0}^{\infty}\frac{(a^{(0)}+b^{(0)}-c_{1}^{(0)}-c_{2}^{(0)}+1)_{k}}{k!}\prod_{i=1}^{2}x_{i}^{k}(1-x_{i})^{c_{i}^{(0)}-a^{(0)}-b^{(0)}-s-k}\,.\end{split}$ (110)
Combining this result with (36) and (104), performing the sum with index $k$, and re-expressing variables in terms of the cross-ratios $(x,z,\bar{z})$, we obtain the leading behavior of the defect-channel conformal blocks in the simultaneous lightcone/large-spin limit:
[…]
up to subleading corrections in the combined limit.